A Bibliography of Publications of Heinz Rutishauser

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

30 January 2021 Version 1.04

Abstract This bibliography records publications of Heinz Rutishauser (30 Jan-

uary 1918–10 November 1970).

Title word cross-reference

1 [Nak66a, Rut66a]. 2 [Nak66b, Rut66b]. $24.50 [G.70]. $3.95 [Str61]. $49 [I.91]. ∫ e^{b(x+a cosx)} dx [Rut51a, Rut51a]. LR [Rut58b, Rut60g, RS63, TT15]. n [Rut55a, Rut60e]. QD [AN68, Rut62c, Ban63a, Dmi13, AC10]. R^{4} [Rut47b].

-algorithm [AC10, Dmi13, Rut62c]. -Algorithmus [Ban63a]. -ter [Rut55a, Rut60e]. -th [Rut55a, Rut60e]. -Transformation [Rut60g, Rut58b].

104 [Rut62a]. 125 [Rut62b]. 135 [Sch69]. 150 [Rut63a, Rut63f]. 1960 [AGG60]. 1969 [Alb71]. 1980 [Ano80].

2 [Rut76b, S.78].

314 [Čul72b]. 34 [H.70]. 34- [H.70]. 367 [Čul72a].

50 [Gep66].

6-ter [HRG59a]. 60 [BBG^{+}60, BBG^{+}63c, BBG^{+}63a, BBG^{+}76, BHO^{+}67, Bec71, Čul72b, Čul72a, Gep66, GHL67, NBB^{+}60, NBB^{+}63, NBB^{+}65, NBB^{+}97, Rut63g, Rut63h, van19]. 6th [HRG59a]. 6x [Rut65].

Abbildungen [Rut50b, Rut50c]. Accelerating [Rut58a]. according [Rut59e]. Adjoint [EGRS59b, Str61]. again [Rut68c]. ago [Gut07]. Akademie [Gep66]. Akademie-Verlag [Gep66]. Al [Rut59b, Rut59b]. Al-Salam [Rut59b, Rut59b]. Algebra [Rut66d, Rut70a, WR71]. algebraic [Rut59d]. ALGOL [BBG^{+}63c, Bec71, Čul72b, Čul72a, Rut13, Gep66, BBG^{+}63c, BBG^{+}63a, BBG^{+}63b, BBG^{+}76, BHO^{+}67, GHL67, NBB^{+}60, NBB^{+}63, NBB^{+}65, Nau68, Rut61, Rut63g, Rut63h, Rut65, Rut13, van19, BBG^{+}60, NBB^{+}97]. ALGOL-60 [BBG^{+}76, NBB^{+}65]. ALGOL-procedure [Rut61]. Algorithm [Ban64, Dmy15, Nak66a, Nak66b, AMM11, AC10, AL98, AN68, Ban60, Ban63a, Ban63b, Dmi13, Dmy16, Rut53, Rut54a, Rut54c, Rut54b, Rut55b, Rut56a, Rut57b, Rut62c, Rut63i, Rut68a, See82, TT15, Rut62a, Rut62b, Rut63a, Rut63f, Rut66a, Rut66b]. algorithme [Rut68a]. Algorithmic [BBG^{+}63c, BBG^{+}76, NBB^{+}60, NBB^{+}63, NBB^{+}97, BBG^{+}60, BBG^{+}63a, BBG^{+}63b, Čul72b, Čul72a, Gep66, NBB^{+}65]. Algorithmische [BBG^{+}63c, Gep66]. Algorithmus [Ban60, Ban63a, Ban63b, Ban64, Nak66a, Nak66b, Rut53, Rut54a, Rut54c, Rut54b, Rut55b, Rut56a, Rut57b, Rut63i, Rut66a, Rut66b, See82]. Algoritm [Rut60f]. Alps [Rut13]. also [Čul72b, Čul72a]. Analogon [Rut54c]. analogue [Rut54c]. Analysis [Par74, SRS73, Gut07]. analytic [Rut50b, Rut50c]. analytical [Rut47b]. analytiques [Rut47b]. analytischen [Rut50b, Rut50c]. Angelitch [G.70]. angewandte [Tod52]. angewandten [Ste69]. Anwendung [Ban63a]. Anwendungen [Rut54a]. Anwendungsgebiete [Sch70, Sch69]. Application [Ban63a]. Applications [Rut67f, Rut54a]. Approach [Fay94, WE90]. approximate [BR68]. Approximation [Rut76a]. Arbeit [Rut51a, Rut59b, Rut13]. aspects [BRS63, Rut69b]. Association [Čul72b]. Aumann [G.70]. Ausdehnung [Rut63b]. Automatic [Bec71, Sch69, Sch70, AN68, Rut51c, Rut57a, Rut52a, Rut52b]. automatico [AN68]. automatische [Rut51c, Rut52a, Rut52b]. automatischen [Rut57a].

B [Čul72b, Čul72a, Gep66, H.70, Bec71, Sch70]. Backus [Čul72b, Čul72a, Gep66]. Band [Bec71, Rut66a, Rut76a, Rut76b, Sch69, Sch70, Rut59c]. Bandmatrix [Rut66a]. Bandmatrizen [Rut59c]. Basel [Str61, Tod52]. Basic [Rut67a]. Bauer [Čul72b, Čul72a, G.70, Gep66, Sch69, Sch70, Rut69b]. Bd [Ste69]. Bedeutung [Rut56a]. Begründung [Ban60]. beim [Rut57a]. Beitrage [Rut53]. Bemerkungen [Rut51a, Rut55a, Rut59b, Rut60d, Rut60e]. Berechnung [HRG59b, Ban57]. Bericht [Rut58d, BBG^{+}63c, Gep66]. Berlin [Bec71, G.70, Gep66]. Bernoulli [AC10]. Berücksichtigung [Sch69, Sch70]. Beschleunigung [Rut58a]. besonderer [Sch69, Sch70]. Besprungen [Bec71]. Bestimmung [Ban63a, Rut58c, Rut55b, Rut66c]. Betrachtungen [Rut63c]. Between [Bau80]. Beweis [See82]. Biografie [Rut13]. Biographies [Wei88]. biography [Rut13]. Biorthogonalisierungs [Rut53]. Biorthogonalisierungs-Algorithmus [Rut53]. biorthogonalization [Rut53]. Birkhauser [I.91, Str61, Tod52]. Book [Alb71, Čul72b, Čul72a, G.70, H.70, I.91, Par74, S.78, Sch69, Sch70, Ste69, Ste78, Str61, Tod52, BSR53, BCK^{+}69, DCR^{+}52, GR58, HHR^{+}68, KPS^{+}67, LWRDR52, LRS^{+}59, LPA^{+}51, MKM^{+}58, MRDJ65, PBJ^{+}52, PRW^{+}56, RDBG^{+}60, RDWZ^{+}62, RFT^{+}57, SBRD^{+}55, SRP52, SGR^{+}67, SSG^{+}63, SRDH^{+}53, WRDG^{+}56, WRDF^{+}63, ZGR62, ZPS^{+}66, vMNJ^{+}51]. Boston [I.91]. Boundary [EGRS59b, Str61]. Bounds [SW81]. Bruhat [TT15]. Buchbesprechung [Gep66]. Buchbesprechungen [BSR53, BCK^{+}69, DCR^{+}52, GR58, HHR^{+}68, KPS^{+}67, LWRDR52, LRS^{+}59, LPA^{+}51, MKM^{+}58, MRDJ65, PBJ^{+}52, PRW^{+}56, RDBG^{+}60, RDWZ^{+}62, RFT^{+}57, SBRD^{+}55, SRP52, SGR^{+}67, SSG^{+}63, SRDH^{+}53, WRDG^{+}56, WRDF^{+}63, ZGR62, ZPS^{+}66, vMNJ^{+}51]. Bulirsch [G.70]. Bulletin [BBG^{+}63c, BBG^{+}63c].

C [Čul72b, Čul72a, Gep66]. calcolo [AN68]. calcul [Bos65, Rut64]. calculating [Rut51c, RSS51b, RSS51c, RSS51a, Rut52a, Rut52b, Rut64]. calculation [AN68, HRG59b, Rut51c, Rut52a, Rut52b, Rut57a, Ban57]. Carlitz [Rut59b]. cases [Rut63i]. Central [Bau80]. Century [MHR80]. Certain [Ban64, Ban63b, Rut58a]. Certification [Nak66a, Nak66b]. chastnykh [Rut60f]. Cholbd [Nak66a]. Cholesky [AL98, Rut66a]. class [Rut58a]. Cliffs [Par74]. closed [RS48]. cm [G.70, H.70, I.91]. cm. [Par74]. Collection [MHR80]. Comment [Rut60a]. Comments [Rut59b]. Communications [Čul72b]. Computation [Bec71, EGRS59b, Sch70, Bos65, Sch69, Str61]. Computational [Rut69b, Alb71]. Computed [SW81]. Computer [Čul72a, Neu05]. Computing [BBR^{+}58, Bau80, Čul72b, MHR80, Sch81, AL98, Rut51b, RSS51b, RSS51c, RSS51a, Sor82, SF88]. Concepts [Rut67a]. concernant [SR61]. Conclusions [EGRS59a]. Considerations [Rut63c]. contains [RS48]. contient [RS48]. Continued [Rut58d, Rut58a]. Contributions [Rut65, Rut53]. controlled [HRG59b, Rut51c, Rut51b, RSS51b, RSS51c, RSS51a, Rut52a, Rut52b]. Convergence [Ban64, Rut50a, RB55, Rut58a, Rut62c]. convergent [Rut60g]. cooperate [AC10]. Correspondence [Rut63d]. courbe [RS48]. cubically [Rut60g]. curve [RS48].

data [GH84]. Declarations [Rut67b]. Definite [Rut66a]. definiter [Rut66a]. Deflation [Rut59d, Rut59c]. Delimiters [Rut60c]. Description [BBR^{+}58, BHO^{+}67, Bec71]. detection [AC10]. Détermination [RB55, Rut55d, Rut58c, Rut55d, Rut55b, RB55, Rut66c]. determine [Ban63a]. determining [Rut69a]. Development [Bau80]. device [Rut51b]. devices [RSS51b, RSS51c, RSS51a]. Difference [Ban64, Ban60, Ban63b, Rut54a, Rut54c, Rut54b, Rut55b, Rut56a, Rut57b, Rut63i, Rut68a, See82]. Differential [Alb71, Rut52c, Rut55a, Rut56b, Rut60e, Rut76b]. Differentialgleichungen [Rut55a, Rut60e, Rut76b, Rut52c]. Differenzen [Ban60, Ban63b, Ban64, Rut54a, Rut54c, Rut55b, Rut56a, See82, Rut54b, Rut57b, Rut63i]. Differenzen-Algorithmus [Ban64, Rut54b, Rut57b, Rut63i]. Digital [Bau80, RSS51b, RSS51c, RSS51a]. digitale [RSS50a, RSS50b, RSS51b, RSS51c, RSS51a, Tod52]. Dimensional [Dmy15, Dmi13, HRG59b]. discussion [Rut65]. distributions [Rut44]. DM [Bec71, H.70]. dont [RS48]. dqds [AMM11]. d’une [RS48, RB55, Rut55d].

E. [Rut66b]. Early [Bau80, Sch81]. Edited [BBG^{+}63c]. editor [Čul72b]. Effort [Nau68]. eigenfunctions [HRG59a]. Eigenfunktionen [HRG59a]. Eigenvalue [Ban64, Fay94, Ban60, Ban63b, Rut58b, Rut59d, Rut76b]. Eigenvalues [EGRS59b, Rut66c, Ban57, Ban63a, Bos65, HRG59b, Rut55b, Rut55d, Rut58c, Rut64, Sor82, SF88, Str61, HRG59a]. eigenvectors [Rut55b, RB55]. Eigenvektoren [Rut55b]. Eigenwerte [Ban57, Ban63a, HRG59b, Rut55b, Rut58c, Rut66c, HRG59a]. Eigenwertprobleme [Ban63b, Ban64, Rut76b]. eindimensionaler [HRG59b]. Einzeldarstellungen [Sch69, Sch70]. electronic [RSS51b, RSS51c, RSS51a]. elektronische [RSS50a, RSS50b, RSS51b, RSS51c, RSS51a, Tod52]. elementary [Rut59d]. Emden [Rut51a]. Engeli [Str61]. Engineer [SS68]. Englewood [Par74]. Equation [Rut66a, Rut76a]. equations [Ban63a, Rut52c, Rut55a, Rut56b, Rut60e, Rut76b, Alb71]. Erfahrungen [Rut51b]. Erinnerung [BBG^{+}63c]. ERMETH [Neu05, Rut55c]. Error [SW81]. Erste [Rut51b]. Erweiterung [Ban63b]. Essays [MHR80]. estate [GMŠ77]. Europe [Bau80]. exact [Zie81b, Zie81a]. exakten [Zie81a]. experiences [Rut51b, Bos65]. experiments [Bos65]. Expressions [Rut60b, Rut67c, Rut60a]. Extension [Ban63b, Rut63b].

F [Čul72b, Čul72a, G.70, Gep66, Sch69, Sch70]. F. [Rut69b, Sch69, Sch70]. factorization [AL98]. Failures [Nau68]. families [Rut47a, Rut47b, Rut50b, Rut50c]. familles [Rut47a, Rut47b]. February [AGG60]. fermée [RS48]. First [Neu05, van19, Rut51b]. Flows [WE90]. Folgen [Rut50b, Rut50c]. fonctions [Rut47a]. Formel [Rut56a]. formula [Rut56a]. fr. [Tod52]. Fractions [Rut58d, Rut58a]. French [Bos65, Rut47a, Rut47b, RS48, RB55, Rut55d, Rut64, Rut68a, SR61]. functions [AN68, HRG59a, Rut47a, Rut50b, Rut50c]. Funktionen [HRG59a, Rut50b, Rut50c]. funzioni [AN68].

G [G.70, H.70]. Gauss [Rut59e, Rut59e]. Gautschi [I.91]. genäherte [BR68]. Generalization [Dmy15, Rut58d, Dmy16]. Generalizations [Zie81b, Dmi13, Zie81a]. generation [AN68]. generazione [AN68]. German [I.91, BBG^{+}63c, Ban57, Ban60, Ban63a, Ban63b, Ban64, BSR53, Bec71, BCK^{+}69, BR68, DCR^{+}52, GMŠ77, GR58, HRG59a, HRG59b, HHR^{+}68, KPS^{+}67, LWRDR52, LRS^{+}59, LPA^{+}51, MKM^{+}58, MRDJ65, Nak66a, Nak66b, PBJ^{+}52, PRW^{+}56, RDBG^{+}60, RDWZ^{+}62, Rut44, Rut50b, Rut50c, Rut50a, Rut51c, Rut51a, Rut51b, RSS51b, RSS51c, RSS51a, Rut52a, Rut52b, Rut52c, Rut53, Rut54a, Rut54c, Rut54b, Rut55b, Rut55a, Rut56a, RFT^{+}57, Rut57b, Rut57a, Rut58c, Rut58a, Rut59b, Rut59c, Rut59e, Rut60d, Rut60e, Rut60g, Rut63b, Rut63c, Rut63i, Rut66a, Rut66b, Rut66c, RSS68, Rut69a, RSS72, Rut76a, Rut76b, Rut13, Rut58d, SS68, SBRD^{+}55, SRP52, See82, SGR^{+}67, SSG^{+}63, SRDH^{+}53, WRDG^{+}56, WRDF^{+}63, ZGR62, Zie81a, ZPS^{+}66, vMNJ^{+}51]. gewissen [Rut58a]. gewisser [Ban63b, Ban64]. gewöhnlicher [Rut52c, Rut55a, Rut60e]. Ginsburg [Str61]. glatten [Rut60d]. Gleichungssystem [Rut66a]. Gleichungssysteme [Rut76a]. Gradient [Rut59a]. Graeffe [Rut62c]. Graeffe-type [Rut62c]. Grau [Bec71]. Green [Čul72b, Čul72a, Gep66]. Grundlehren [Bec71, Sch69, Sch70]. Gutknecht [S.78].

H [Alb71, Bec71, Čul72b, Čul72a, G.70, Gep66, H.70, Par74, Sch69, Sch70, Ste69, Str61, Tod52]. H. [H.70, Par74, Ste69]. Hadamard [AC10]. Hall [Par74]. Handbook [Rut66d, Rut70a, Bec71, Sch69, Sch70]. Hans [S.78]. Hans-Rudolf [S.78]. Heidelberg [Bec71]. Heinz [Anoxx, GMŠ77, Rut13, Ste78, Ano10, Bau71, Bau05, GMŠ77, I.91, OR10, Rut13, S.78]. help [Rut55b]. Henrici [S.78]. Herausgegeben [BBG^{+}63c]. Hertelendy [Par74]. Hg [Sch70]. High [Rut13]. Hilfe [Rut55b]. Hilfsmittel [G.70, SS68]. Hill [Bec71]. History [MHR80]. Hochalpen [Rut13]. Householder [Sch69, Sch70].

ibid [Čul72a]. ihre [Rut56a]. II [RSS50b]. III [G.70, RSS51b, SS68]. improvement [Rut50a]. infinitesimal [Rut54c]. infinitesimales [Rut54c]. Ingenieurs [G.70, SS68]. Input [Rut67d]. Insights [van19]. Instabilität [Rut52c]. instability [Rut52c, Rut56b]. Institut [Tod52]. integral [Ban63a]. Integralgleichungen [Ban63a]. integrating [Rut52c]. Integration [Rut52c, Rut55a, Rut60e, Rut56b, SR61, SR61]. Interference [Rut61]. Interpolation [Rut76a, Rut60d, BR68, Rut60d]. Introduction [Rut67e]. Invariant [SW81]. Inversen [Zie81a]. inverses [Zie81b, Zie81a]. inversion [Rut59e]. Italian [AN68]. iteration [Rut63c, Rut69b, Rut70a, Rut70b, Rut71b, TT15]. Iterative [EGRS59b, RB55, Str61, RB55]. IV [RSS51c].

J [Čul72b, Čul72a, G.70, Gep66, Par74, Sch69, Sch70]. J. [Čul72b, Čul72a]. Jacobi [GH84, Rut62a, Rut63e, Rut66d, Rut66e, Rut71a]. Jordan [Rut59e]. Journal [Čul72a]. justification [Ban60].

K. [Rut51a, Rut51b]. Katz [Čul72b, Čul72a, Gep66]. Kentniss [Rut53]. Kerner [Gep66]. Kettenbrüche [Rut58d]. Kettenbrüchen [Rut58a]. Klasse [Rut58a]. konvergente [Rut60g]. Konvergenz [Ban64, Rut58a]. Konvergenzverbesserung [Rut50a]. kubisch [Rut60g]. Kugelfläche [Rut44]. Kunzi [G.70].

L [Čul72b, Čul72a, G.70, Gep66, Rut69b, Sch69, Sch70]. L’algoritmo [AN68]. Lanczos [Rut53]. Langmaack [Bec71]. Language [BBG^{+}63c, BBG^{+}76, BBR^{+}58, NBB^{+}60, NBB^{+}63, NBB^{+}97, BBG^{+}60, BBG^{+}63a, BBG^{+}63b, Čul72b, Čul72a, Gep66, NBB^{+}65]. Lauchli [S.78]. least [Rut68c]. Lectures [Rut90, I.91, Rut76a, Rut76b]. Lehre [GMŠ77]. Leitfaden [Ste69]. Letters [AGG60]. Linear [Ban64, Rut66a, Rut66d, Rut70a, WR71, Ban63b, HRG59a]. linearen [HRG59a]. linearer [Ban63b, Ban64]. Lineares [Rut66a]. Lösung [Rut51a].

M [Čul72b, Čul72a, Gep66, Str61]. Machinery [Čul72b]. machines [Rut51c, RSS51b, RSS51c, RSS51a, Rut52a, Rut52b]. maps [Rut50b, Rut50c]. Martin [S.78]. mathematical [SS68]. Mathematics [Rut90, Rut76a, Rut76b, I.91]. Mathematik [Čul72b, Čul72a, Rut76a, Rut76b, S.78, Tod52, Ste69, Ste78]. Mathematische [G.70, SS68]. mathematischen [Bec71, Sch69, Sch70]. matrice [RB55, Rut55d]. Matrices [Rut68b, Rut71a, SRS73, GH84, Rut58c, Rut59c, RS63, Rut64, Rut66d, Rut66e, RSS68, Rut70a, Rut70b, Rut71b, RSS72, Sor82, SF88, Par74, Rut66c]. Matrix [Ban57, Rut55b, Rut66a, RSS68, Ban60, RB55, Rut55d, Rut59e, Zie81b, Zie81a]. Matrixeigenwertprobleme [Ban60]. Matrizen [RSS68, Ste69, H.70, Rut58c, Rut66c, RSS68, RSS72, Alb71]. Matrizen-Numerik [RSS68]. Matrizeninversion [Rut59e]. McCarthy [Čul72b, Čul72a, Gep66]. MDN [Gep66]. meaning [Rut56a]. Mechanik [Ste69]. mehrerer [Rut50b, Rut50c]. memory [BBG^{+}63c]. meromorphen [Rut50b, Rut50c]. méromorphes [Rut47a]. meromorphic [Rut50b, Rut50c]. Method [Rut71a, Ban57, Bos65, Rut50a, RB55, Rut55d, RS63, Rut64, Rut66d, Rut66e, Rut69b, Rut70a, Rut70b, Rut71b, Sor82, SF88]. méthode [Bos65, RB55, Rut55d, Rut64, Rut50a]. Methoden [Rut52c]. Methods [Alb71, EGRS59b, Rut59a, Rut52c, Rut56b, Str61]. metomorphic [Rut47a]. Miscellaneous [Rut67f]. Mitarb [Gep66]. Mitchell [Alb71]. Mitteilungen [Tod52]. Modelle [HRG59b]. models [HRG59b]. modification [Rut62c]. Modified [BBG^{+}76, Sor82, SF88]. molecular [HRG59b]. Molekel [HRG59b]. Molekel-Modelle [HRG59b]. Moore [Zie81a, Zie81b, Zie81a]. multidimensional [Dmy16]. My [Bau02, Bau03, Bau11].

N [Par74]. nach [Rut59e, Rut66a, Rut66b]. Nachlass [GMŠ77]. Naur [Čul72b, Čul72a, Gep66, AGG60, BBG^{+}63c]. nel [AN68]. Newton [Rut50a]. Newtonsche [Rut50a]. No [BBG^{+}63c, Čul72a, Tod52]. non [Rut64]. note [AMM11, Kas57]. Notes [Rut60d, SR61]. Nullstellenbestimmung [Rut69a]. Numerical [Gut07, I.91, Rut68a, Rut90, SRS73, BRS63, Bos65, Rut55a, Rut60e, Rut76a, Rut76b, SR61, GMŠ77, Par74]. numerically [GH84]. numerics [RSS68, RSS68, RSS72, Rut13]. Numerik [Ste69, H.70, RSS68, RSS72, Rut13, Alb71]. numérique [SR61]. numériques [Bos65, Rut68a]. Numerische [Čul72b, Čul72a, S.78, Rut76a, Rut76b, Ste78, GMŠ77]. numerischen [Rut55a, Rut60e].

O [Gep66]. October [Ano80]. Oh [Wei88]. Old [van19]. Oliver [Sch70]. Olver [Sch69]. Once [Rut68c]. one [HRG59b]. one-dimensional [HRG59b]. Operatoren [Ban63b, Ban64]. Operators [Ban64, Ban63b]. order [Rut55a, Rut60e]. ordinary [Rut52c, Rut55a, Rut56b, Rut60e]. Ordnung [Rut55a, Rut60e]. ornach [Nak66b]. Orthogonal [Rut66c]. orthogonaler [Rut66c]. Orthonormalization [Rut66b]. Orthonormierung [Rut66b]. oscillator [HRG59a]. Oszillators [HRG59a]. Output [Rut67d]. overrelaxation [EGRS59c].

P [BBG^{+}63c, Čul72b, Čul72a, G.70, Gep66, Par74]. P. [BBG^{+}63c]. Part [Bec71, Sch70, Rut76a, SS68]. Partial [Alb71]. patterns [Rut63e]. Penrose [Zie81b, Zie81a]. Penrose-Inversen [Zie81a]. Perlis [Čul72b, Čul72a, Gep66]. Peter [S.78, AGG60]. Pioneers [Wei88]. plan [Rut51c, Rut52a, Rut52b]. plusieurs [Rut47a]. point [Rut44]. Polynomen [Rut69a]. polynomials [Rut69a]. positiv [Rut66a]. positiv-definiter [Rut66a]. Positive [Rut66a]. Potenz [HRG59a]. pour [Bos65, Rut55d, Rut64]. power [HRG59a]. pp [Čul72b, Čul72a, G.70, H.70, I.91, Par74, Str61, Tod52]. Preis [Gep66]. Preliminary [AGG60, Rut58d]. Prentice [Par74]. Prentice-Hall [Par74]. preparatory [Rut57a]. Price [G.70, H.70, I.91]. principle [Rut63b]. Prinzips [Rut63b]. problem [Rut68c, Rut69a]. Problematik [Rut69a]. Problems [Ban64, EGRS59b, Fay94, Ban60, Ban63b, Rut58b, Rut59d, Rut76b, Str61]. procedure [Rut61, HRG59b]. Procedures [Rut63h, Rut67g, GMŠ77, Rut63g]. Processes [BBR^{+}58]. production [Rut51c, Rut52a, Rut52b]. program [HRG59b, Rut51c, Rut51b, Rut52a, Rut52b, RSS51b, RSS51c, RSS51a]. program-controlled [HRG59b, Rut51c, Rut51b, Rut52a, Rut52b, RSS51b, RSS51c, RSS51a].

Programmgesteuerte [RSS50a, RSS50b, RSS51b, RSS51c, RSS51a, Tod52]. programmgesteuerten [HRG59b, Rut51c, Rut51b, Rut52a, Rut52b]. Programming [Rut55c]. Programs [van19]. proof [See82]. properties [Rut68a]. Proposal [BBR^{+}58]. propres [Bos65, RB55, Rut55d, Rut64]. propriétés [Rut68a]. Prozeduren [GMŠ77]. Punktverteilungen [Rut44].

qd [Dmy15, Dmy16]. qd-Algorithm [Dmy15, Dmy16]. quadratic [RB55]. quadratique [RB55]. Quadratur [BR68]. quadrature [BRS63, BR68]. Quadratwurzeliteration [Rut63c]. Quelques [Bos65]. Quotient [Ban64, Ban60, Ban63b, Rut54a, Rut54c, Rut54b, Rut55b, Rut56a, Rut57b, Rut63i, Rut68a, See82]. quotient-difference [Ban60]. Quotienten [Ban60, Ban63b, Ban64, Rut54b, Rut57b, Rut63i, Rut54a, Rut54c, Rut55b, Rut56a, See82]. Quotienten-Differenzen-Algorithmus [Ban60, Ban63b, Rut54a, Rut54c, Rut55b, Rut56a, See82].

R [Alb71, G.70, H.70, Par74, Ste69]. R. [G.70]. radius [RS48]. rayon [RS48]. raznostei [Rut60f]. Real [Rut71a, Ban57, Rut66d, Rut66e, SF88]. Rechengerät [Rut51b]. Rechengeräte [RSS50a, RSS50b, RSS51b, RSS51c, RSS51a, Tod52]. Rechenmaschinen [RSS50a, RSS50b, RSS51b, RSS51c, RSS51a, Tod52, Rut51c, Rut52a, Rut52b]. Rechenplanfertigung [Rut51c, Rut52a, Rut52b]. Rechnen [Rut57a]. reconstruction [GH84]. Recursive [Rut63h, Rut63g]. Reduction [Rut62a]. reellen [Ban57]. Refined [EGRS59b, Str61]. Regeln [See82]. relations [Rut60a, Rut60b]. Reliable [AC10]. Remark [Rut60c, Rut63f]. Remarks [Rut51a, Rut55a, Rut60e]. Remarques [SR61]. Report [AGG60, BBG^{+}60, BBG^{+}76, NBB^{+}60, NBB^{+}63, NBB^{+}97, Rut58d, BBG^{+}63a, BBG^{+}63b, Čul72a, NBB^{+}65, BBG^{+}63c, Čul72b, Gep66]. representations [Rut47b, Rut47b]. Review [Čul72b, Čul72a, G.70, H.70, I.91, Par74, S.78, Sch69, Sch70, Ste69, Ste78, Str61, Tod52]. reviews [Alb71, BSR53, BCK^{+}69, DCR^{+}52, GR58, HHR^{+}68, KPS^{+}67, LWRDR52, LRS^{+}59, LPA^{+}51, MKM^{+}58, MRDJ65, PBJ^{+}52, PRW^{+}56, RDBG^{+}60, RDWZ^{+}62, RFT^{+}57, SBRD^{+}55, SRP52, SGR^{+}67, SSG^{+}63, SRDH^{+}53, WRDG^{+}56, WRDF^{+}63, ZGR62, ZPS^{+}66, vMNJ^{+}51, Bec71]. Revised [BBG^{+}63a, BBG^{+}63b, NBB^{+}63, NBB^{+}65, NBB^{+}97, Čul72a]. Romberg [Rut63b]. Rombergschen [Rut63b]. root [AC10, Rut63c]. rotation [Rut63e]. rotations [Rut59d]. Rudolf [S.78]. rule [See82]. Russian [NBB^{+}65]. Ruthishauser [Tod52]. Rutishauser [Alb71, Ban60, Ban63a, Ban63b, Ban64, Bec71, Čul72b, Čul72a, G.70, Gep66, H.70, I.91, Par74, S.78, Sch69, Sch70, Ste69, Str61, Zie81a, AMM11, AC10, AL98, Ano80, Ano10, Anoxx, AN68, Ban57, Ban60, Ban63a, Ban63b, Ban64, Bau71, Bau80, Bau02, Bau03, Bau05, Bau11, Bos65, Dmi13, Dmy15, Dmy16, Fay94, GMŠ77, Kas57, OR10, Rut13, See82, Sor82, SF88, Ste78, TT15, WE90, Zie81b, Zie81a].

S [Bec71, Gep66, Sch69, Sch70]. Salam [Rut59b, Rut59b]. Samelson [Čul72b, Čul72a, G.70, Gep66, Sch69, Sch70]. Sauer [G.70]. Scharen [Rut50b, Rut50c]. schiefsymmetrischer [Rut58c]. Schmidt [Rut66b]. Schwarz [Alb71, H.70, Par74, S.78, Ste69]. Schweizer [Rut13]. Self [EGRS59b, WE90, Str61]. Self-Adjoint [EGRS59b, Str61]. Self-Similar [WE90]. sequences [Rut47a, Rut47b, Rut50b, Rut50c]. Series [Rut66d, Rut70a]. several [Rut47a, Rut50b, Rut50c]. shift [AMM11, AL98]. Similar [WE90]. Simultaneous [Rut70a, Rut70b, Rut71b, Rut69b]. singular [AMM11]. skew [Rut58c]. skew-symmetric [Rut58c]. smooth [Rut60d]. Solution [EGRS59b, Rut58b, Rut51a, Rut59d, Str61]. Some [Rut55c, Bos65]. Sonderfälle [Rut63i]. sowie [Rut50b, Rut50c]. special [AN68, Rut63i]. speciali [AN68]. spectral [GH84]. Speiser [Tod52]. sphere [RS48, RS48]. spherical [Rut44]. Sprache [BBG^{+}63c, Gep66]. Springer [G.70, Bec71]. Springer-Verlag [G.70, Bec71]. square [Rut63c, Rut68c]. Stabile [Rut63i]. stable [GH84, Rut63i]. Statements [Rut67h]. Stiefel [Alb71, H.70, Par74, Sch69, Sch70, Ste69, Str61, Tod52]. Stoer [G.70]. strategy [AL98]. Structure [van19]. Stuttgart [H.70, Str61]. Subspaces [SW81]. Successes [Nau68]. suites [Rut47a, Rut47b]. Superfluous [Rut60c]. Supplement [BBG^{+}63c]. surface [Rut44, RS48]. SVD [AL98]. Swiss [Tod52, Neu05, Rut13]. Switzerland [Sch81]. symétriques [Rut64]. Syminv2 [Rut63a, Rut63f]. Symmetric [Par74, Rut66a, Rut71a, SRS73, Rut58c, RS63, Rut66d, Rut66e, RSS68, Rut70a, Rut70b, Rut71b, RSS72, Sor82, SF88]. Symmetrischer [Alb71, H.70, Rut66a, RSS68, RSS72, Ste69]. Symposium [Ano80]. System [van19]. Systems [Rut66a, Rut76a]. Szabo [G.70].

T [G.70]. teaching [GMŠ77]. Techniques [Rut55c]. Teil [SS68]. ter [HRG59a, Rut55a, Rut60e]. Test [Rut68b, Zie81b, Zie81a]. Testmatrix [Zie81a]. Teubner [H.70]. Th [Str61, Rut55a, Rut60e]. theorem [Kas57]. Theory [Rut59a]. thermodynamic [HRG59a]. thermodynamische [HRG59a]. tools [SS68]. transformation [Rut58b, Rut60g, RS63, Rut60g]. translated [I.91, Par74]. Translation [Bec71, GHL67]. tridiagonal [SF88]. Turanski [BBG^{+}63c, BBG^{+}63c]. Twentieth [MHR80]. Two [Dmi13, Dmy15]. Two-Dimensional [Dmy15, Dmi13]. type [Rut62c].

U [Bec71, Gep66]. Übers [Gep66]. understanding [Rut53]. Universal [BBR^{+}58]. unsymmetrical [Rut64]. Use [Rut63h, Rut63g]. using [Ban57, Rut66a].

valeurs [Bos65, Rut55d, Rut64]. Value [EGRS59b, Str61]. values [AMM11]. Variabeln [Rut50b, Rut50c]. variables [Rut47a, Rut50b, Rut50c]. variant [Rut60g]. Variante [Rut60g]. Vauquois [Čul72b, Čul72a, Gep66]. vecteurs [RB55]. Vectors [Rut66b]. Vektoren [Rut66b]. Verallgemeinerung [Rut58d]. Verallgemeinerungen [Zie81a]. Verfahren [Ban57, HRG59b]. Verlag [G.70, Gep66, H.70, Str61, Bec71]. vi [Gep66]. Vol [Bec71, Sch69, Sch70, Čul72b, Čul72a]. Vols [S.78]. Volume [Rut76b]. Vorbereitungsarbeit [Rut57a]. Vorläufiger [Rut58d]. Vorlesungen [S.78, Rut76a, Rut76b, Ste78].

W [Čul72b, Čul72a, Gep66, Sch70]. Walter [I.91]. Wegstein [Čul72b, Čul72a, Gep66]. Weightcoeff [Rut62b]. well [Rut50b, Rut50c]. whose [RS48]. Wijngaarden [Čul72b, Čul72a]. Wijngarden [Gep66]. William [BBG^{+}63c, BBG^{+}63c]. Wissenschaften [Bec71, Sch69, Sch70]. Woodger [Čul72b, Čul72a, Gep66]. work [Rut51a, Rut57a, Rut59b, Rut13]. Wronski [Rut56a, Rut56a].

X [Bec71]. xi [Par74]. XII [Bec71]. xv [I.91]. xviv [G.70].

year [Bau03]. Years [Sch81, Bau02, Bau11, Gut07]. York [Bec71]. Young [EGRS59c].

Z4 [Rut51b, Rut51b]. zeros [Rut69a]. Zertifikat [Nak66a, Nak66b]. zum [Rut54c]. Zur [BBG^{+}63c, Rut13, Ban63a, HRG59b, Rut51a, Rut52c, Rut53, Rut55a, Rut60d, Rut60e, Rut63c, Ban60, Rut58c, Rut59e, Rut69a]. Zurich [Ano80, Gut07]. Zuse [Bau80, Rut51b].

References

Allouche:2010:RRD

[AC10] Hassane Allouche and Annie Cuyt. Reliable root detection with the qd-algorithm: When Bernoulli, Hadamard and Rutishauser cooperate. Applied Numerical Mathematics: Transactions of IMACS, 60(12):1188–1208, December 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Arden:1960:LPR

[AGG60] B. Arden, B. Galler, and R. Graham. Letters: Preliminary report (February 4, 1960) of Peter Naur. Communications of the ACM, 3 (6):A13, June 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [NBB^{+}60].

Andrisani:1998:SCS

[AL98] E. Andrisani and G. Di Lena. A strategy for computing the shift in the Cholesky–Rutishauser algorithm for the SVD factorization. Rendiconti di matematica e delle sue applicazioni (1981), 18(4): 725–748 (1999), 1998. CODEN RNMTAN. ISSN 1120-7183.

Albaseny:1971:BRN

[Alb71] E. L. Albaseny. Book reviews: Numerik Symmetrischer Matrizen, by H. Rutishauser, E. Stiefel and H. Schwarz, 1968; Computational Methods in Partial Differential Equations, by A. R. Mitchell, 1969. The Computer Journal, 14(4):374, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/tiff/374.tif.

Aishima:2011:NDA

[AMM11] Kensuke Aishima, Takayasu Matsuo, and Kazuo Murota. A note on the dqds algorithm with Rutishauser’s shift for singular values. Japan Journal of Industrial and Applied Mathematics, 28(2):251– 262, ???? 2011. CODEN JAPJI7. ISSN 0916-7005 (print), 1868-

937x (electronic). URL http://link.springer.com/article/
10.1007/s13160-011-0037-x.

Ascari:1968:LRG

[AN68] A. Ascari and P. G. Novario. L’algoritmo QD di Rutishauser e la generazione di funzioni speciali nel calcolo automatico. (Italian) [The QD algorithm of Rutishauser and the generation of special functions in automatic calculation]. Calcolo: a quarterly on numerical analysis and theory of computation, 5(1):162–173, ???? 1968. CODEN CALOBK. ISSN 0008-0624 (print), 1126-

5434 (electronic). URL http://link.springer.com/article/
10.1007/BF02576063.

Anonymous:1980:RSZ

[Ano80] Anonymous. Rutishauser Symposium, Zurich, October 15–17, 1980. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 31(2):??, ???? 1980. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01590756.

Anonymous:2010:HR

[Ano10] Anonymous. Heinz Rutishauser. Historisches Lexikon der Schweiz (HLS) Web site, July 12, 2010. URL https://hls-dhs-dss.ch/

de/articles/031649/2010-07-12/.

Anonymous:20xx:RH

[Anoxx] Anonymous. Rutishauser, Heinz (1918–1970). Swiss elites

Web site., 20xx. URL https://www2.unil.ch/elitessuisses/
index.php?page=detailPerso&idIdentite=77302.

Bandemer:1956:BRE

[Ban57] Hans Bandemer. Berechnung der reellen Eigenwerte einer reellen Matrix mit dem Verfahren von Rutishauser. (German) [Calculation of the real eigenvalues of a real matrix using the Rutishauser method]. Wiss. Z. Martin-Luther-Univ. Halle-Wittenberg Math.- Natur. Reihe, 6:807–814, 1956–1957. ISSN 0138-1504.

Bandemer:1960:BQD

[Ban60] H. Bandemer. Zur Begründung des Quotienten-Differenzen-Algorithmus von Rutishauser für Matrixeigenwertprobleme. (German) [On the justification for Rutishauser’s quotient-difference algorithm for matrix eigenvalue problems]. Wiss. Z. Martin-Luther-Univ. Halle-Wittenberg Math.-Natur. Reihe, 9:263–268, 1960. ISSN 0138-1504.

Bandemer:1963:AAR

[Ban63a] Hans Bandemer. Anwendung des QD-Algorithmus von Rutishauser zur Bestimmung der Eigenwerte von Integralgleichungen. (German) [Application of the QD algorithm by Rutishauser to determine the eigenvalues of integral equations]. Wiss. Z. Techn. Univ. Dresden, 12:112–114, 1963. ISSN 0043-6925.

Bandemer:1963:EQD

[Ban63b] Hans Bandemer. Erweiterung des Quotienten-Differenzen-Algorithmus von Rutishauser auf Eigenwertprobleme gewisser linearer Operatoren. (German) [Extension of the quotient–difference algorithm of Rutishauser to eigenvalue problems of certain linear operators]. Wiss. Z. Martin-Luther-Univ. Halle-Wittenberg Math.-Natur. Reihe, 12:235–249, 1963. ISSN 0138-1504.

Bandemer:1963:KQD

[Ban64] Hans Bandemer. Über die Konvergenz des Quotienten– Differenzen-Algorithmus von Rutishauser für Eigenwertprobleme gewisser linearer Operatoren. (German) [On the convergence of Rutishauser’s quotient–difference algorithm for eigenvalue problems of certain linear operators]. Mathematische Nachrichten, 27: 353–375, 1963–1964. CODEN MTMNAQ. ISSN 0025-584X.

Bauer:1971:HR

[Bau71] F. L. Bauer. Heinz Rutishauser (10.11.1970). Computing, 7(1– 2):129–130, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https://link.springer.

com/article/10.1007/BF02279949.

Bauer:1980:BZR

[Bau80] Friedrich L. Bauer. Between Zuse and Rutishauser — the early development of digital computing in Central Europe. In Metropolis et al. [MHR80], pages 505–524. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976.

URL http://www.sciencedirect.com/science/article/pii/

B978012491650050037X. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Bauer:2002:MYR

[Bau02] Friedrich L. Bauer. My years with Rutishauser. In ????, editor, LATSIS Symposium 2002: ETH, Zürich, Switzerland,

page ?? ????, ????, 2002. URL http://www.cs.umd.edu/users/
oleary/cggg/bauer.pdf;   https://vimeo.com/308173562/
5b0d69f62f.

Bauer:2003:MYR

[Bau03] Friedrich L. Bauer. My year with Rutishauser. DINAL, 11(1–4): 1–13, 2003.

Bauer:2005:RH

[Bau05] Friedrich L. Bauer. Rutishauser, Heinz. In Neue Deutsche Biographie (NDB). (German) [New German Biography], page 305. Duncker & Humblot, Berlin, Germany, 2005. URL https:/

/daten.digitale-sammlungen.de/0001/bsb00016410/images/
index.html?seite=319.

Bauer:2011:MYR

[Bau11] Friedrich L. Bauer. My years with Rutishauser. Informatik Spektrum, 34(5):514–518, October 2011. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://link.

springer.com/article/10.1007/s00287-011-0554-7.

Backus:1960:RAL

[BBG^{+}60] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Report on the algorithmic language Algol 60. Numerische Mathematik, 2(1): 106–136, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/BF01386216.

Backus:1963:RRAa

[BBG^{+}63a] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Revised report on the algorithmic language AL-GOL 60. The Computer Journal, 5(4):349–367, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/5/
4/349.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/050349.sgm.abs.html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_04/tiff/349.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/350.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
04/tiff/351.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/352.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/353.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/354.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/355.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
04/tiff/356.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/357.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/358.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/359.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/360.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
04/tiff/361.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/362.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/363.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/364.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/365.tif;   http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
04/tiff/366.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/367.tif.

Backus:1963:RRAb

[BBG^{+}63b] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Revised report on the algorithmic language ALGOL 60. Numerische Mathematik, 4:420–453, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/BF01386340. Edited by Peter

Naur. Dedicated to the memory of William Turanski.

Backus:1963:BAS

[BBG^{+}63c] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, P. Naur, A. J. Perlis, Heinz Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Bericht über die Algorithmische Sprache ALGOL 60 (ALGOL Bulletin Supplement No. 3). Zur Erinnerung an William Turanski. Herausgegeben von P. Naur. (German) [Report on the algorithmic language ALGOL 60 (ALGOL Bulletin supplement no. 3). In memory of William Turanski. Edited by P. Naur]. Elektronische Datenverarbeitung, 2(??):1–17, 1963. ISSN 0374-3012.

Backus:1976:MRA

[BBG^{+}76] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, M. Woodger, R. M. De Morgan, I. D. Hill, and B. A. Wichmann. Modified report on the algorithmic language ALGOL-60. The Computer Journal, 19(4):364–379, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/4/364.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_04/tiff/364.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/
Issue_04/tiff/365.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_04/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
04/tiff/367.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/368.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_19/Issue_04/tiff/369.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_04/tiff/370.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_04/tiff/371.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
04/tiff/372.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/373.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_19/Issue_04/tiff/374.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_04/tiff/375.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_04/tiff/376.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
04/tiff/377.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/378.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_19/Issue_04/tiff/379.
tif.

Bauer:1958:PUL

[BBR^{+}58] F. L. Bauer, H. Bottenbruch, H. Rutishauser, K. Samelson, P. Graeff, P. Lauchli, and M. Paul. Proposal for a universal language for the description of computing processes. In J. W. Carr, editor, Computer Programming and Artificial Intelligence, University of Michigan Summer School 1958, pages 355–373. ????, ????, 1958.

Buhlmann:1969:BGB

[BCK^{+}69] H. Bühlmann, W. Czaja, K. Kovari, R. T. Shield, W. Bosshard, K. Alder, and H. Rutishauser. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 20(3):430– 432, May 1969. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-

9039 (electronic). URL http://link.springer.com/article/
10.1007/BF01590436.

Becker:1971:BRH

[Bec71] Otwin Becker. Besprungen: Rutishauser, H.: Handbook for Automatic Computation, Vol. I / Part a: Description of ALGOL 60, Die Grundlehren der mathematischen Wissenschaften, 135. Band. XII, 323 S. DM 58,-. Grau, A. A., U. Hill, H. Langmaack: Handbook for Automatic Computation, Vol. 1 / Part b: Translation of ALGOL 60, Die Grundlehren der mathematischen Wissenschaften, 137. Band. X, 397 S. DM 64,-. Berlin– Heidelberg–New York 1967, Springer-Verlag. (German) [Reviews]. Zeitschrift für die gesamte Staatswissenschaft / Journal of Institutional and Theoretical Economics, 127(2):365–366, May 1971.

CODEN ???? ISSN 0044-2550. URL https://www.jstor.org/

stable/40749460.

Bauer:1967:DA

[BHO^{+}67] Friedrich L. Bauer, Alston S. Householder, Frank W. J. Olver, Heinz Rutishauser, Klaus Samelson, and Eduard Stiefel. Description of ALGOL 60, volume 1a of Handbook for Automatic Computation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. ISBN 3-540-03826-4, 3-642-86934-3 (e-book), 3-642-86936-X (print), 3-662-38103-6. xii + 326

pp. LCCN QA76.5 .R87. URL http://www.springerlink.com/
content/978-3-642-86934-1.

Bosset:1965:QEN

[Bos65] L. Bosset. Quelques expériences numériques sur la méthode de Rutishauser pour le calcul des valeurs propres. (French) [Some numerical experiments on the method of Rutishauser for the computation of eigenvalues]. In Troisième Congr. de Calcul et de Traitement de l’Information AFCALTI, pages 123–129. Dunod, Paris, France, 1965.

Bulirsch:1968:IGQ

[BR68] R. Bulirsch and H. Rutishauser. Interpolation und genäherte Quadratur. (German) [Interpolation and approximate quadrature]. In Sauer and Szabó [SS68], pages 232–319. ISBN 3-642-95030-2, 3-642-95031-0. LCCN QA76 .B3192. Herausgegeben von R. Sauer und I. Szabó. Unter Mitwirkung von H. Neuber, W. Nürnberg, K. Pöschl, E. Truckenbrodt und W. Zander.

Bauer:1963:NAN

[BRS63] F. L. Bauer, H. Rutishauser, and E. Stiefel. New aspects in numerical quadrature. In Proceedings of the Symposium on Applied Mathematics, volume XV, pages 199–218. American Mathematical Society, Providence, RI, USA, 1963.

Baumgartner:1953:BGB

[BSR53] W. Baumgartner, E. Stiefel, and H. Rutishauser. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 4(1):87–88, January 1953. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http://link.springer.

com/article/10.1007/BF02075310.

Culik:1972:BRJb

[Čul72a] Karel Čulík. Book review: J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Revised report on the algorithmic language ALGOL 60. Numerische Mathematik, vol. 4 (1962–1963), pp. 420–453; also ibid., vol. 6 (1963), pp. 1–17; and The Computer Journal, vol. 5 no. 4 (1963), pp. 349–367. Journal of Symbolic Logic, 37(3):623– 624, September 1972. CODEN JSYLA6. ISSN 0022-4812 (print),

1943-5886 (electronic). URL https://www.jstor.org/stable/
2272794.

Culik:1972:BRJa

[Čul72b] Karel Čulík. Book review: J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur (editor), A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Report on the algorithmic language ALGOL 60. Numerische Mathematik, vol. 2 (1960), pp. 106–136; also Communications of the Association for Computing Machinery, vol. 3 (1960), pp. 299–314. Journal of Symbolic Logic, 37(3):622–623, September 1972. CODEN JSYLA6. ISSN 0022-4812 (print),

1943-5886 (electronic). URL ttps://www.jstor.org/stable/
2272793.

Dunner:1952:BGB

[DCR^{+}52] E. Dünner, B. Chaix, H. Rutishauser, M. Plancherel, and H. Ziegler. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 3(4):317–320, July 1952. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02009625.

Dmitrishin:2013:TDG

[Dmi13] R. Ī. Dmitrishin. Two-dimensional generalizations of the Rutishauser qd-algorithm. Mat. Metodi Fiz.-Mekh. Polya, 56(4): 33–39, 2013. ISSN 0130-9420.

Dmytryshyn:2015:TDG

[Dmy15] R. I. Dmytryshyn. Two-dimensional generalization of the Rutishauser qd-algorithm. Journal of mathematical sciences (New York, N.Y.), 208(3):301–309, June 2015. CODEN JMTSEW.

ISSN 1072-3374 (print), 1573-8795 (electronic). URL http://

link.springer.com/article/10.1007/s10958-015-2447-9.

Dmytryshyn:2016:MGR

[Dmy16] R. I. Dmytryshyn. A multidimensional generalization of the Rutishauser qd-algorithm. Carpathian Math. Publ., 8(2):230–238, 2016. ISSN 2075-9827.

Engeli:1959:C

[EGRS59a] M. Engeli, Th. Ginsburg, H. Rutishauser, and E. Stiefel. Conclusions. In Refined Iterative Methods for Computation of the Solution and the Eigenvalues of Self-adjoint Boundary Value Problems [EGRS59b], chapter V, pages 102–105. ISBN 3-0348-7224-0 (e-book), 3-0348-7226-7. ISSN 0514-8146. LCCN QA3 .Z8 v.8;

QA371. URL http://www.springerlink.com/content/978-3-
0348-7224-9.

Engeli:1959:RIM

[EGRS59b] Max Engeli, Theo. Ginsburg, Heinz Rutishauser, and Eduard L. Stiefel. Refined Iterative Methods for Computation of the Solution and the Eigenvalues of Self-Adjoint Boundary Value Problems, volume 8 of Mitteilungen aus dem Institut für angewandte Mathematik. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1959. ISBN 3-0348-7224-0 (e-book), 3-0348-7226-7. ISSN 0514-8146. 107 pp. LCCN QA3 .Z8 v.8;

QA371. URL http://www.springerlink.com/content/978-3-
0348-7224-9.

Engeli:1959:YO

[EGRS59c] Max Engeli, Theo. Ginsburg, Heinz Rutishauser, and Eduard L. Stiefel. Young’s overrelaxation. In Refined Iterative Methods for Computation of the Solution and the Eigenvalues of Self-adjoint Boundary Value Problems [EGRS59b], chapter IV.1, page ?? ISBN 3-0348-7224-0 (e-book), 3-0348-7226-7. ISSN 0514-8146.

LCCN QA3 .Z8 v.8; QA371. URL http://www.springerlink.
com/content/978-3-0348-7224-9.

Faybusovich:1994:RAE

[Fay94] L. Faybusovich. On the Rutishauser’s approach to eigenvalue problems. In Paul Van Dooren and Bostwick Wyman, editors, Linear Algebra for Control Theory, volume 62 of The IMA Volumes in Mathematics and its Applications, pages 87–102. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London,

UK / etc., 1994. URL http://link.springer.com/chapter/
10.1007/978-1-4613-8419-9_8.

G:1970:BRR

[G.70] W. G. Book review: R. Sauer, I. Szabo, T. P. Angelitch, G. Aumann, F. L. Bauer, R. Bulirsch, H. P. Kunzi, H. Rutishauser, K. Samelson, R. Sauer, and J. Stoer, Mathematische Hilfsmittel des Ingenieurs. III, Springer-Verlag, Berlin, 1969, xviv + 534 pp., 24 cm. Price $24.50. Mathematics of Computation, 24(110): 475, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print),

1088-6842 (electronic). URL https://www.jstor.org/stable/
2004494.

Geppert:1966:BBJ

[Gep66] M. G. Geppert. Buchbesprechung: Backus, J. W., F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. Van Wijngarden und M. Woodger: Bericht über die algorithmische Sprache Algol 60. (Report on the algorithmic language Algol 60. Übers. Kerner J. O. u. Mitarb.). Akademie-Verlag Berlin 1962. vi + 49 S., Preis MDN 10,50. Biometrische Zeitschrift, 8(1-2):121, 1966. CODEN BIZEB3. ISSN 0006-3452.

Gragg:1984:NSR

[GH84] William B. Gragg and William J. Harrod. The numerically stable reconstruction of Jacobi matrices from spectral data. Numerische Mathematik, 44(3):317–335, September 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Grau:1967:TA

[GHL67] A. A. Grau, U. Hill, and H. Langmaack. Translation of AL-GOL 60, volume 137 of Handbook for Automatic Computation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. ISBN 3-642-86937-8 (e-book), 3-642-86939-4 (print). ISSN 0072-7830 (print), 2196-9701 (electronic). ix +

403 pp. LCCN QA1-939. URL http://www.springerlink.com/
content/978-3-642-86937-2.

Gander:1977:NPN

[GMŠ77] Walter Gander, Luciano Molinari, and Hana Švecová, editors. Numerische Prozeduren aus Nachlass und Lehre von Heinz Rutishauser. (German) [Numerical procedures from the estate and teaching of Heinz Rutishauser], volume 33 of International Series of Numerical Mathematics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1977. ISBN 3-7643-0874-5, 3-0348-7176-7 (e-book). 127 pp. LCCN QA297 .R86 1977.

Gruter:1958:BGB

[GR58] H. Grüter and H. Rutishauser. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 9(4): 393–394, November 1958. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http://link.springer.

com/article/10.1007/BF01593881.

Gutknecht:2007:NAZ

[Gut07] Martin H. Gutknecht. Numerical analysis in Zurich — 50 years ago. In ????, editor, Proceedings of ICIAM 2007, pages 279–290. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN ???? LCCN

???? URL http://www.sam.math.ethz.ch/~mhg/SAMhist/pub/
Gut10-NAZ-SMG100.pdf.

H:1970:BRH

[H.70] A. S. H. Book review: H. R. Schwarz, H. Rutishauser & E. Stiefel, Numerik Symmetrischer Matrizen, B. G. Teubner Verlag, Stuttgart, 1968, 243 pp., 22 cm. Price DM 34-. Mathematics of Computation, 24(109):229–230, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

https://www.jstor.org/stable/2004900.

Henrich:1968:BGB

[HHR^{+}68] P. Henrich, P. Huguenin, H. Rutishauser, I. Zschokke-Gränacher, Hans Ziegler, H. R. Schwarz, L. S. Dzung, and E. Stiefel. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 19(6):981–984, November 1968. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http://

link.springer.com/article/10.1007/BF01602290.

Heilbronner:1959:EET

[HRG59a] E. Heilbronner, H. Rutishauser, and F. Gerson. Eigenwerte, Eigenfunktionen und thermodynamische Funktionen des linearen Oszillators 6-ter Potenz. (German) [Eigenvalues, eigenfunctions and thermodynamic functions of the linear oscillator of 6th power].

Helvetica Chimica Acta, 42(7):2304–2314, 1959. CODEN HCA-

CAV. ISSN 0018-019X. URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/hlca.19590420703.

Heilbronner:1959:VPB

[HRG59b] E. Heilbronner, H. Rutishauser, and F. Gerson. Verfahren zur programmgesteuerten Berechnung der Eigenwerte eindimensionaler Molekel-Modelle. (German) [Procedure for the program-controlled calculation of the eigenvalues of one-dimensional molecular models]. Helvetica Chimica Acta, 42(7):2285–2303, 1959. CODEN HCACAV. ISSN 0018-019X. URL https://onlinelibrary.

wiley.com/doi/pdf/10.1002/hlca.19590420702.

I:1991:BRH

[I.91] E. I. Book review: Heinz Rutishauser, Lectures on Numerical Mathematics (translated from the German by Walter Gautschi), Birkhauser, Boston, 1990, xv + 546 pp., 23.5 cm. Price $49. Mathematics of Computation, 57(196):869–870, October 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

https://www.jstor.org/stable/2938724.

Kasajima:1957:NTR

[Kas57] Tomomi Kasajima. A note on a theorem of Rutishauser. Commentarii Mathematici Universitatis Sancti Pauli, 6:89–91, 1957. ISSN 0010-258X.

Kuhn:1967:BGB

[KPS^{+}67] W. Kuhn, P. Profos, H. R. Schwarz, Theo Ginsburg, H. Rutishauser, K. Alder, Ch. Wehrli, and Peter J. Huber. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 18 (1):145–148, January 1967. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http://link.springer.

com/article/10.1007/BF01593907.

Linder:1951:BGB

[LPA^{+}51] A. Linder, M. Plancherel, J. Ackeret, H. J. v. Baeyer, A. A. Rusterholz, A. Linder, H. Rutishauser, H. Ziegler, F. Kobold, and F. Stüssi. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 2(1):53–62, January 1951. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02589817.

Lauchli:1959:BGB

[LRS^{+}59] P. Läuchli, H. Rutishauser, R. Sänger, H. Weber, E. Roth-Desmeules, C. Enz, Hs. H. Günthard, V. Stingelin, W. Hälg, F. Heinrich, and H. Ziegler. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 10(4):444–452, July 1959. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-

9039 (electronic). URL http://link.springer.com/article/
10.1007/BF01601058.

Labhart:1952:BGB

[LWRDR52] H. Labhart, H. Wäffler, E. Roth-Desmeules, and H. Rutishauser. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 3(6):477–478, November 1952. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL

http://link.springer.com/article/10.1007/BF02025579.

Metropolis:1980:HCT

[MHR80] Nicholas Metropolis, Jack Howlett, and Gian-Carlo Rota, editors. A History of Computing in the Twentieth Century: a Collection of Essays. Academic Press, New York, NY, USA, 1980. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Maeder:1958:BGB

[MKM^{+}58] D. Maeder, U. Katz, E. Mooser, P. de Haller, W. Saxer, H., R. Voellmy, R. Jost, H. Rutishauser, and H. Ziegler. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 9(2):211–216, July 1958. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http://link.

springer.com/article/10.1007/BF01600636.

McVean:1965:BGB

[MRDJ65] D. B. McVean, H. Rutishauser, L. S. Dzung, and M. Jeger. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 16(4):566–568, July 1965. CODEN ZAM-PDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http:

//link.springer.com/article/10.1007/BF01593957.

Nake:1966:ZAC

[Nak66a] F. Nake. Zertifikat zu Algorithmus 1, Cholbd. (German) [Certification of algorithm 1, Cholbd]. Computing, 1(1):79, March 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [Rut66a].

Nake:1966:ZAO

[Nak66b] F. Nake. Zertifikat zu Algorithmus 2 ornach. (German) [Certification of algorithm 2]. Computing, 1(3):281, September 1966. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [Rut66b].

Naur:1968:SFA

[Nau68] Peter Naur. Successes and failures of the ALGOL effort. AL-GOL Bulletin (Amsterdam: Mathematisch Centrum), ??(28):58– 62, July 1968. CODEN ALGOBG. ISSN 0084-6198. URL

http://doi.acm.org/10.1145/1061112.1061118. Talk given at

the colloquium “Ten Years of Algol”, Eidgenössische Technische Hochschule, Zürich, May 31, 1968.

Naur:1960:RALa

[NBB^{+}60] Peter Naur, J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Report on the algorithmic language ALGOL 60. Communications of the ACM, 3(5):299–314, May 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See letter [AGG60]. Dedicated to the Memory of William Turanski [American group member who was killed by an automobile just prior to the January 1960 Conference].

Naur:1963:RRA

[NBB^{+}63] Peter Naur, J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, M. Woodger, and P. Nauer. Revised report on the algorithmic language ALGOL 60. Communications of the ACM, 6(1):1–17, January 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Dedicated to the memory of William Turanski.

Naur:1965:RRA

[NBB^{+}65] P. Naur, J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, P. Naur, A. J. Perlis, Heinz Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Revised report on the algorithmic language ALGOL-

60. (Russian), volume Moskau, Russia. Verlag “Mir”, 1965. 80 pp.

Naur:1997:RRA

[NBB^{+}97] Peter Naur, J. W. Backus, F. L. Bauer, J. Green, C. Kafz, J. Mc-Carthy, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Revised report on the algorithmic language Algol 60. In Algol-like Languages, pages 19–49. Birkhäuser Boston Inc., Cambridge, MA, USA, 1997.

Neukom:2005:EFS

[Neu05] Hans Neukom. ERMETH: The first Swiss computer. IEEE Annals of the History of Computing, 27(4):5–22, October/December 2005. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

OConnor:2010:HR

[OR10] John J. O’Connor and Edward F. Robertson. Heinz Rutishauser. MacTutor History of Mathematics Web site, St. Andrews University, Scotland., December 2010. URL https://mathshistory.

st-andrews.ac.uk/Biographies/Rutishauser/;   https://
www.wikiwand.com/en/MacTutor_History_of_Mathematics_archive.

Parlett:1974:BRH

[Par74] Beresford N. Parlett. Book review: H. R. Schwarz, H. Rutishauser & E. Stiefel, translated by P. Hertelendy, Numerical Analysis of Symmetric Matrices, Prentice-Hall, Inc., Englewood Cliffs, N. J., 1973, xi + 276 pp., 24 cm. Mathematics of Computation, 28(125): 328, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print),

1088-6842 (electronic). URL https://www.jstor.org/stable/
2005845.

Plancherel:1952:BGB

[PBJ^{+}52] M. Plancherel, W. Baumgartner, M. Jeger, H. Rutishauser, A. P. Speiser, J. Ackeret, and E. Völlm. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 3(5): 401–406, September 1952. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http://link.springer.

com/article/10.1007/BF02008156.

Plancherel:1956:BGB

[PRW^{+}56] M. Plancherel, H. Rutishauser, H. Weber, J. Ackeret, and H. Ziegler. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 7(1):86–88, January 1956. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01600732.

Rutishauser:1955:DVP

[RB55] Heinz Rutishauser and Friedrich L. Bauer. Détermination des vecteurs propres d’une matrice par une méthode itérative avec convergence quadratique. (French) [Determination of the eigenvectors of a matrix by an iterative method with quadratic convergence]. Comptes rendus de l’Académie des sciences, Paris, 240: 1680–1681, 1955. ISSN 0001-4036 (print), 2419-6304 (electronic).

Roth-Desmeules:1960:BGB

[RDBG^{+}60] E. Roth-Desmeules, H. J. Bürgin, Th. Ginsburg, W. Hardmeier, Res Jost, H. Ziegler, F. Tank, W. Traupel, and H. Rutishauser. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 11(1):81–88, January 1960. CODEN ZAM-PDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http:

//link.springer.com/article/10.1007/BF01591809.

Roth-Desmeules:1962:BGB

[RDWZ^{+}62] E. Roth-Desmeules, H. Weber, H. Ziegler, E. Baldinger, H. Rutishauser, M. Weber, J.-P. Blaser, and Res Jost. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 13(6):613–620, November 1962. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http://

link.springer.com/article/10.1007/BF01595588.

Rutishauser:1957:BGB

[RFT^{+}57] H. Rutishauser, H. Favre, F. Tank, W. Hälg, E. Roth-Desmeules, B., L. van der Waerden, P. Profos, P. de Haller, and E. Bickel.

Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 8(1):82–88, January 1957. CODEN ZAM-PDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http:

//link.springer.com/article/10.1007/BF01601159.

Rutishauser:1948:RSD

[RS48] Heinz Rutishauser and Hans Samelson. Sur le rayon d’une sphère dont la surface contient une courbe fermée. (French) [On the radius of a sphere whose surface contains a closed curve]. Comptes rendus de l’Académie des sciences, Paris, 227:755–757, 1948. ISSN 0001-4036 (print), 2419-6304 (electronic).

Rutishauser:1963:TMS

[RS63] H. Rutishauser and H. R. Schwarz. The LR transformation method for symmetric matrices. Numerische Mathematik, 5(1): 273–289, December 1963. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic).

Rutishauser:1950:PDRa

[RSS50a] Heinz Rutishauser, Ambros Speiser, and Eduard Stiefel. Programmgesteuerte digitale Rechengeräte (elektronische Rechenmaschinen). I. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 1(5):277–297, September 1950. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Rutishauser:1950:PDRb

[RSS50b] Heinz Rutishauser, Ambros Speiser, and Eduard Stiefel. Programmgesteuerte digitale Rechengeräte (elektronische Rechenmaschinen). II. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 1(5):339–362, September 1950. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Rutishauser:1951:PDRc

[RSS51a] Heinz Rutishauser, Ambros Speiser, and Eduard Stiefel. Programmgesteuerte digitale Rechengeräte (elektronische Rechenmaschinen). (German) [Program-controlled digital computing devices (electronic calculating machines)], volume 2 of Mitteilungen aus dem Institut für angewandte Mathematik. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1951. 102 pp.

URL https://link.springer.com/article/10.1007/978-3-
0348-4113-9;   https://link.springer.com/book/10.1007/
978-3-0348-4113-9.

Rutishauser:1951:PDRa

[RSS51b] Heinz Rutishauser, Ambros Speiser, and Eduard Stiefel. Programmgesteuerte digitale Rechengeräte (elektronische Rechenmaschinen). III. (German) [Program-controlled digital computing devices (electronic calculating machines). III]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 2(1):1–25, January 1951. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Rutishauser:1951:PDRb

[RSS51c] Heinz Rutishauser, Ambros Speiser, and Eduard Stiefel. Programmgesteuerte digitale Rechengeräte (elektronische Rechenmaschinen). IV. (German) [Program-controlled digital computing devices (electronic calculating machines). IV]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 2(2):62–92, March 1951. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Rutishauser:1968:NSM

[RSS68] Heinz Rutishauser, Eduard L. Stiefel, and Hans Rudolf Schwarz. Numerik symmetrischer Matrizen: Matrizen-Numerik. (German) [Numerics of symmetric matrices: Matrix numerics], volume 11 of Leitfäden der angewandten Mathematik und Mechanik. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1968. ISSN 0459-021X. 243 pp. LCCN QA263 .S33.

Rutishauser:1972:NSM

[RSS72] H. Rutishauser, E. Stiefel, and H. R. Schwarz. Numerik symmetrischer Matrizen. (German) [Numerics of symmetric matrices], volume 11 of Leitfäden der angewandten Mathematik und Mechanik LAMM. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1972. ISBN 3-519-12311-8 (print), 3-663-11341-8 (ebook). 261 pp. LCCN QA263 .S33 1972.

Rutishauser:1944:PKG

[Rut44] H. Rutishauser. Über Punktverteilungen auf der Kugelfläche. (German) [On point distributions on the spherical surface]. Commentarii Mathematici Helvetici (Zurich), 17(1):327–331, December 1944. CODEN COMHAX. ISSN 0010-2571 (print), 1420-8946 (electronic).

Rutishauser:1947:SFF

[Rut47a] H. Rutishauser. Sur les suites et familles de fonctions méromorphes de plusieurs variables. (French) [On the sequences and families of metomorphic functions of several variables]. Comptes rendus de l’Académie des sciences, Paris, 224:1804–1806, 1947. ISSN 0001-4036 (print), 2419-6304 (electronic).

Rutishauser:1947:SFR

[Rut47b] Heinz Rutishauser. Sur les suites et familles de représentations analytiques du R^{4}. (French) [On sequences and families of analytical representations of R^{4}]. Comptes rendus de l’Académie des sciences, Paris, 225:33–35, 1947. ISSN 0001-4036 (print), 2419-6304 (electronic).

Rutishauser:1950:KNM

[Rut50a] Heinz Rutishauser. Eine Konvergenzverbesserung für die Newtonsche Methode. (German) [A convergence improvement for Newton’s method]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 1:211–212, 1950. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Rutishauser:1950:FSAa

[Rut50b] Heinz Rutishauser. Über Folgen und Scharen von analytischen und meromorphen Funktionen mehrerer Variabeln, sowie von analytischen Abbildungen. (German) [On sequences and families of analytic and meromorphic functions of several variables, as well as of analytic maps]. Ph.D. thesis, Eidgenössische Technische Hochschule (ETH), Zurich, Switzerland, 1950. ii + 249–325 pp.

Rutishauser:1950:FSAb

[Rut50c] Heinz Rutishauser. Über Folgen und Scharen von analytischen und meromorphen Funktionen mehrerer Variabeln, sowie von analytischen Abbildungen. (German) [On sequences and families of analytic and meromorphic functions of several variables, as well as of analytic maps]. Acta Mathematica, 83:249–325, 1950. CO-DEN ACMAA8. ISSN 0001-5962 (print), 1871-2509 (electronic). This is the author’s Ph.D. thesis at ETH.

Rutishauser:1951:BAK

[Rut51a] Heinz Rutishauser. Bemerkungen zur Arbeit von K. Emden “Eine Lösung für ∫ e^{b(x+a cos x)} dx”. . (German) [Remarks on the work by K. Emden, “A solution for ∫ e^{b(x+a cosx)} dx”]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 2(4):292–293, July 1951. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Rutishauser:1951:EEP

[Rut51b] Heinz Rutishauser. Erste Erfahrungen mit dem programmgesteuerten Rechengerät Z4 von K. Zuse. (German) [First experiences with the program-controlled computing device Z4 from K. Zuse]. Zeitschrift für Angewandte Mathematik und Mechanik, 31 (8–9):254, 1951. CODEN ZAMMAX. ISSN 0044-2267 (print),

1521-4001 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/zamm.19510310819.

Rutishauser:1951:ARP

[Rut51c] Heinz Rutishauser. Über automatische Rechenplanfertigung bei programmgesteuerten Rechenmaschinen. (German) [On automatic calculation plan production with program-controlled calculating machines]. Zeitschrift für Angewandte Mathematik und Mechanik, 31(8–9):255, 1951. CODEN ZAM-MAX. ISSN 0044-2267 (print), 1521-4001 (electronic). URL

https://onlinelibrary.wiley.com/doi/abs/10.1002/zamm.
19510310820.

Rutishauser:1952:ARPa

[Rut52a] Heinz Rutishauser. Automatische Rechenplanfertigung bei programmgesteuerten Rechenmaschinen. (German) [Automatic calculation plan production with program-controlled calculating machines]. Mitt. Inst. Angew. Math. Zürich, 1952(3):45, 1952. ISSN 0514-8146.

Rutishauser:1952:ARPb

[Rut52b] Heinz Rutishauser. Automatische Rechenplanfertigung bei programmgesteuerten Rechenmaschinen. (German) [Automatic calculation plan production with program-controlled calculating machines]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 3(4):312–313, July 1952. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Rutishauser:1952:IMI

[Rut52c] Heinz Rutishauser. Über die Instabilität von Methoden zur Integration gewöhnlicher Differentialgleichungen. (German) [On the instability of methods for integrating ordinary differential equations]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 3(1):65–74, January 1952. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). See English translation in [Rut56b].

Rutishauser:1953:BKB

[Rut53] Heinz Rutishauser. Beitrage zur Kentniss des Biorthogonalisierungs-Algorithmus von Lanczos. (German) [Contributions to the understanding of the Lanczos biorthogonalization algorithm]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 4(1):35–56, January 1953. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Rutishauser:1954:AQD

[Rut54a] Heinz Rutishauser. Anwendungen des Quotienten-Differenzen-Algorithmus. (German) [Applications of the quotient–difference algorithm]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 5(6):496–508, November 1954. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Rutishauser:1954:QDA

[Rut54b] Heinz Rutishauser. Der Quotienten–Differenzen-Algorithmus. (German) [The quotient–difference algorithm]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 5(3):233–251, May 1954. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Rutishauser:1954:IAQ

[Rut54c] Heinz Rutishauser. Ein infinitesimales Analogon zum Quotienten-Differenzen-Algorithmus. (German) [An infinitesimal analogue to the quotient–difference algorithm]. Archiv der Mathematik, 5:132– 137, 1954. CODEN ACVMAL. ISSN 0003-889x (print), 1420-8938 (electronic).

Rutishauser:1955:BNI

[Rut55a] Heinz Rutishauser. Bemerkungen zur numerischen Integration gewöhnlicher Differentialgleichungen n-ter Ordnung. (German)

[Remarks on the numerical integration of ordinary differential equations n-th order]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 6(6):497– 498, November 1955. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Rutishauser:1955:BEE

[Rut55b] Heinz Rutishauser. Bestimmung der Eigenwerte und Eigenvektoren einer Matrix mit Hilfe des Quotienten-Differenzen-Algorithmus. (German) [Determination of the eigenvalues and eigenvectors of a matrix with the help of the quotient–difference algorithm]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 6(5):387–401, September 1955. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Rutishauser:1955:SPT

[Rut55c] Heinz Rutishauser. Some programming techniques for the ER-METH. Journal of the ACM, 2(1):1–4, January 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735x (electronic). URL

https://dl.acm.org/doi/10.1145/320789.320790.

Rutishauser:1955:MPD

[Rut55d] Heinz Rutishauser. Une méthode pour la détermination des valeurs propres d’une matrice. (French) [A method for the determination of the eigenvalues of a matrix]. Comptes rendus de l’Académie des sciences, Paris, 240:34–36, 1955. ISSN 0001-4036 (print), 2419-6304 (electronic).

Rutishauser:1956:FWI

[Rut56a] Heinz Rutishauser. Eine Formel von Wronski und ihre Bedeutung für den Quotienten-Differenzen-Algorithmus. (German) [A formula by Wronski and its meaning for the quotient–difference algorithm]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 7(2):164–169, March 1956. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Rutishauser:1956:IMI

[Rut56b] Heinz Rutishauser. On the instability of methods for the integration of ordinary differential equations. NACA Technical Memorandum, 1956(1403):15, April 1956. ISSN 0096-7602. URL https://

archive.org/details/nasa_techdoc_20030065241;   https:

//digital.library.unt.edu/ark:/67531/metadc63924/;   https:
//ufdc.ufl.edu/AA00009205/00001. Translation of [Rut52c].

Rutishauser:1957:VBA

[Rut57a] H. Rutishauser. Über die Vorbereitungsarbeit beim automatischen Rechnen. (German) [On the preparatory work for automatic calculation]. Mitt. Verein. Schweiz. Versich.-Math., 57:247–257, 1957. ISSN 0042-3815.

Rutishauser:1957:QDA

[Rut57b] Heinz Rutishauser. Der Quotienten–Differenzen-Algorithmus. (German) [The quotient–difference algorithm]. Mitt. Inst. Angew. Math. Zürich, 1957(7):74, 1957. ISSN 0514-8146.

Rutishauser:1958:BKG

[Rut58a] Heinz Rutishauser. Beschleunigung der Konvergenz einer gewissen Klasse von Kettenbrüchen. (German) [Accelerating the convergence of a certain class of continued fractions]. Zeitschrift für Angewandte Mathematik und Mechanik, 38(5–6):187–190, May/ June 1958. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-

4001 (electronic). URL https://onlinelibrary.wiley.com/
doi/abs/10.1002/zamm.19580380506.

Rutishauser:1958:SEP

[Rut58b] Heinz Rutishauser. Solution of eigenvalue problems with the LRtransformation. In Further Contributions to the Solution of Simultaneous Linear Equations and the Determination of Eigenvalues, volume 1958(49) of Applied Mathematics Series, pages 47–81. National Bureau of Standards, United States Department of Commerce, Washington, DC, USA, 1958. ISSN 0083-1786.

Rutishauser:1958:BES

[Rut58c] Heinz Rutishauser. Zur Bestimmung der Eigenwerte schiefsymmetrischer Matrizen. (German) [On determination of the eigenvalues of skew-symmetric matrices]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 9(5–6):586–590, March 1958. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Ruttishauser:1958:VKV

[Rut58d] Heinz Ruttishauser. Über eine Verallgemeinerung der Kettenbrüche (Vorläufiger Bericht). (German) [On a generalization of continued fractions (preliminary report)]. Zeitschrift für Angewandte Mathematik und Mechanik, 38(7–8):278–279, 1958. CO-DEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.
1002/zamm.19580380716.

Rutishauser:1959:TGM

[Rut59a] H. Rutishauser. Theory of gradient methods. In Refined Iterative Methods for Computation of the Solution and the Eigenvalues of Self-adjoint Boundary Value Problems [EGRS59b], chapter II, pages 24–49. ISBN 3-0348-7224-0 (e-book), 3-0348-7226-7. ISSN 0514-8146. LCCN QA3 .Z8 v.8; QA371. URL http://www.

springerlink.com/content/978-3-0348-7224-9.

Rutishauser:1959:BAA

[Rut59b] Heinz Rutishauser. Bemerkungen zu einer Arbeit von Al-Salam und Carlitz. (German) [Comments on a work by Al-Salam and Carlitz]. Archiv der Mathematik, 10:292–293, 1959. CODEN ACVMAL. ISSN 0003-889x (print), 1420-8938 (electronic).

Rutishauser:1959:DBG

[Rut59c] Heinz Rutishauser. Deflation bei Bandmatrizen. (German) [Deflation in band matrices]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 10(3): 314–319, May 1959. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Rutishauser:1959:DER

[Rut59d] Heinz Rutishauser. Deflation by elementary rotations for the solution of algebraic eigenvalue problems. In ACM’59: Preprints of papers presented at the 14th national meeting of the Association for Computing Machinery, September 1959, pages 1–3. ACM Press, New York, NY 10036, USA, 1959.

Rutishauser:1959:MNG

[Rut59e] Heinz Rutishauser. Zur Matrizeninversion nach Gauss–Jordan. (German) [On matrix inversion according to Gauss–Jordan]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 10(3):281–291, May 1959. CO-DEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Rutishauser:1960:CER

[Rut60a] H. Rutishauser. Comment on expressions in relations. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), ??(11):7, December 1960. CODEN ALGOBG. ISSN 0084-6198. URL https://

dl.acm.org/doi/10.5555/1060905.1060912.

Rutishauser:1960:ER

[Rut60b] H. Rutishauser. Expressions in relations. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), ??(10):11, October 1960. CO-DEN ALGOBG. ISSN 0084-6198. URL https://dl.acm.org/

doi/10.5555/1060892.1060896.

Rutishauser:1960:RSD

[Rut60c] H. Rutishauser. Remark on superfluous delimiters. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), ??(11):6, December 1960. CODEN ALGOBG. ISSN 0084-6198.

Rutishauser:1960:BGI

[Rut60d] Heinz Rutishauser. Bemerkungen zur glatten Interpolation. (German) [Notes on smooth interpolation]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 11(6):508–513, November 1960. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Rutishauser:1960:BNI

[Rut60e] Heinz Rutishauser. Bemerkungen zur numerischen Integration gewöhnlicher Differentialgleichungen n-ter Ordnung. (German) [Remarks on the numerical integration of ordinary differential equations of n-th order]. Numerische Mathematik, 2(1):263–279, December 1960. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic).

Rutishauser:1960:CAC

[Rut60f] Heinz Rutishauser. Algoritm chastnykh i raznosteĭ. Biblioteka Sbornika “Matematika”. Izdat. Inostr. Lit., Moscow, USSR, 1960. 93 pp. Translated from the German by V. M. Kuročkin.

Rutishauser:1960:KKV

[Rut60g] Heinz Rutishauser. Über eine kubisch konvergente Variante der LR-Transformation. (German) [On a cubically convergent variant of the LR transformation]. Zeitschrift für Angewandte Mathematik und Mechanik, 40(1–3):49–54, ???? 1960. CO-DEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.
1002/zamm.19600400109.

Rutishauser:1961:IAP

[Rut61] H. Rutishauser. Interference with an ALGOL-procedure. Annual Review in Automatic Programming, 2:67–75, January 1961. CO-DEN ARAPDX. ISSN 0066-4138 (print), 1878-545X (electronic).

Rutishauser:1962:ARJ

[Rut62a] H. Rutishauser. Algorithm 104: Reduction to Jacobi. Communications of the ACM, 5(7):387, July 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rutishauser:1962:AW

[Rut62b] H. Rutishauser. Algorithm 125: Weightcoeff. Communications of the ACM, 5(10):510–511, October 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/368959.368977.

Rutishauser:1962:MAG

[Rut62c] Heinz Rutishauser. On a modification of the QD-algorithm with Graeffe-type convergence. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 13(5): 493–496, September 1962. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Rutishauser:1963:AS

[Rut63a] H. Rutishauser. Algorithm 150: Syminv2. Communications of the ACM, 6(2):67–68, February 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rutishauser:1963:ARP

[Rut63b] H. Rutishauser. Ausdehnung des Rombergschen Prinzips. (German) [Extension of Romberg’s principle]. Numerische Mathematik, 5(1):48–54, December 1963. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic).

Rutishauser:1963:BQG

[Rut63c] H. Rutishauser. Betrachtungen zur Quadratwurzeliteration. (German) [Considerations on square root iteration]. Monatshefte für Mathematik, 67:452–464, 1963. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081 (electronic).

Rutishauser:1963:C

[Rut63d] H. Rutishauser. Correspondence. The Computer Journal, 6(2): 133, August 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/2/133.full.pdf+html.

Rutishauser:1963:JRP

[Rut63e] H. Rutishauser. On Jacobi rotation patterns. In Proc. Sympos. Appl. Math., Vol. XV, pages 219–239. American Mathematical Society, Providence, RI, USA, 1963.

Rutishauser:1963:RAS

[Rut63f] H. Rutishauser. Remark on Algorithm 150: Syminv2. Communications of the ACM, 6(7):390, July 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rutishauser:1963:URPa

[Rut63g] H. Rutishauser. The use of recursive procedures in ALGOL 60. Annual Review in Automatic Programming, 3:43–51, January 1963. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (electronic).

Rutishauser:1963:URPb

[Rut63h] H. Rutishauser. The use of recursive procedures in ALGOL 60. In International Tracts in Computer Science and Technology and Their Application, pages 43–51. Elsevier, Amsterdam, The Netherlands, 1963.

Rutishauser:1963:SSQ

[Rut63i] Heinz Rutishauser. Stabile Sonderfälle des Quotienten–Differenzen-Algorithmus. (German) [Stable special cases of the quotient– difference algorithm]. Numerische Mathematik, 5(1):95–112, December 1963. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic).

Rutishauser:1964:MPC

[Rut64] Heinz Rutishauser. Une méthode pour le calcul des valeurs propres des matrices non symétriques. (French) [A method for calculating the eigenvalues of unsymmetrical matrices]. Comptes rendus de l’Académie des sciences, Paris, 259:2758, 1964. ISSN 0001-4036 (print), 2419-6304 (electronic).

Rutishauser:1965:CDA

[Rut65] H. Rutishauser. Contributions to the discussion of ALGOL 6x. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), ??(19): 44–50, January 1965. CODEN ALGOBG. ISSN 0084-6198. URL

https://dl.acm.org/doi/10.5555/1060998.1061010.

Rutishauser:1966:ALG

[Rut66a] H. Rutishauser. Algorithmus 1: Lineares Gleichungssystem mit symmetrischer positiv-definiter Bandmatrix nach Cholesky. (German) [Algorithm 1: Linear equation systems with symmetric positive definite band matrix using Cholesky]. Computing, 1(1):77–78, March 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-

5057 (electronic). URL http://link.springer.com/article/

10.1007/BF02235854. See certification [Nak66a].

Rutishauser:1966:AOV

[Rut66b] H. Rutishauser. Algorithmus 2: Orthonormierung von Vektoren nach E. Schmidt. (German) [Algorithm 2: Orthonormalization of vectors by E. Schmidt]. Computing, 1(2):159–161, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF02342626. See certification [Nak66b].

Rutishauser:1966:BEO

[Rut66c] H. Rutishauser. Bestimmung der Eigenwerte orthogonaler Matrizen. (German) [Determination of the Eigenvalues of Orthogonal Matrices]. Numerische Mathematik, 9(2):104–108, December 1966. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic).

Rutishauser:1966:HSL

[Rut66d] H. Rutishauser. Handbook series linear algebra: The Jacobi method for real symmetric matrices. Numerische Mathematik, 9(1):1–10, 1966. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic).

Rutishauser:1966:JMR

[Rut66e] H. Rutishauser. The Jacobi method for real symmetric matrices. Numerische Mathematik, 9(1):1–10, November 1966. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 202–211].

Rutishauser:1967:BC

[Rut67a] H. Rutishauser. Basic concepts. In Description of AL-GOL 60 [BHO^{+}67], pages 18–33. ISBN 3-540-03826-4, 3-642-86934-3 (e-book), 3-642-86936-X (print), 3-662-38103-6. LCCN

QA76.5 .R87. URL http://link.springer.com/chapter/10.
1007/978-3-642-86934-1_2.

Rutishauser:1967:D

[Rut67b] H. Rutishauser. Declarations. In Description of ALGOL 60 [BHO^{+}67], pages 144–171. ISBN 3-540-03826-4, 3-642-86934-3 (e-book), 3-642-86936-X (print), 3-662-38103-6. LCCN

QA76.5 .R87. URL http://link.springer.com/chapter/10.
1007/978-3-642-86934-1_6.

Rutishauser:1967:E

[Rut67c] H. Rutishauser. Expressions. In Description of ALGOL 60 [BHO^{+}67], pages 34–55. ISBN 3-540-03826-4, 3-642-86934-3 (ebook), 3-642-86936-X (print), 3-662-38103-6. LCCN QA76.5

.R87. URL http://link.springer.com/chapter/10.1007/
978-3-642-86934-1_3.

Rutishauser:1967:IO

[Rut67d] H. Rutishauser. Input and output. In Description of AL-GOL 60 [BHO^{+}67], pages 227–247. ISBN 3-540-03826-4, 3-642-86934-3 (e-book), 3-642-86936-X (print), 3-662-38103-6. LCCN

QA76.5 .R87. URL http://link.springer.com/chapter/10.
1007/978-3-642-86934-1_8.

Rutishauser:1967:I

[Rut67e] H. Rutishauser. Introduction. In Description of ALGOL 60 [BHO^{+}67], pages 1–17. ISBN 3-540-03826-4, 3-642-86934-3 (ebook), 3-642-86936-X (print), 3-662-38103-6. LCCN QA76.5

.R87. URL http://link.springer.com/chapter/10.1007/
978-3-642-86934-1_1.

Rutishauser:1967:MA

[Rut67f] H. Rutishauser. Miscellaneous applications. In Description of ALGOL 60 [BHO^{+}67], pages 95–143. ISBN 3-540-03826-4, 3-642-86934-3 (e-book), 3-642-86936-X (print), 3-662-38103-6. LCCN

QA76.5 .R87. URL http://link.springer.com/chapter/10.
1007/978-3-642-86934-1_5.

Rutishauser:1967:P

[Rut67g] H. Rutishauser. Procedures. In Description of ALGOL 60 [BHO^{+}67], pages 172–226. ISBN 3-540-03826-4, 3-642-86934-3 (e-book), 3-642-86936-X (print), 3-662-38103-6. LCCN QA76.5

.R87. URL http://link.springer.com/chapter/10.1007/
978-3-642-86934-1_7.

Rutishauser:1967:S

[Rut67h] H. Rutishauser. Statements. In Description of ALGOL 60 [BHO^{+}67], pages 56–94. ISBN 3-540-03826-4, 3-642-86934-3 (ebook), 3-642-86936-X (print), 3-662-38103-6. LCCN QA76.5

.R87. URL http://link.springer.com/chapter/10.1007/
978-3-642-86934-1_4.

Rutishauser:1968:PNA

[Rut68a] H. Rutishauser. Les propriétés numériques de l’algorithme quotient–difference. (French) [numerical properties of the quotient– difference algorithm]. Report EUR 4083f, Communauté Européenne de l’Énergie Atomique — EURATOM, Luxembourg, 1968.

Rutishauser:1968:TM

[Rut68b] H. Rutishauser. On test matrices. In Programmation en Mathématiques Numériques, Besançon, 1966, volume 7(165) of Éditions Centre Nat. Recherche Sci., Paris, pages 349–365. Éditions Centre Nat. Recherche Sci., Paris, France, 1968.

Rutishauser:1968:OAL

[Rut68c] H. Rutishauser. Once again: The least square problem. Linear Algebra and its Applications, 1(4):479–488, October 1968. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0024379568900220.

Rutishauser:1969:PNP

[Rut69a] H. Rutishauser. Zur Problematik der Nullstellenbestimmung bei Polynomen. (German) [On the problem of determining zeros in polynomials]. In Constructive Aspects of the Fundamental Theorem of Algebra (Proc. Sympos., Zürich-Rüschlikon, 1967), pages 281–294. Wiley-Interscience, New York, NY, USA, 1969.

Rutishauser:1969:CAF

[Rut69b] Heinz Rutishauser. Computational aspects of F. L. Bauer’s simultaneous iteration method. Numerische Mathematik, 13(1):4–13, March 1969. CODEN NUMMA7. ISSN 0029-599x (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/BF02165269.

Rutishauser:1970:HSL

[Rut70a] H. Rutishauser. Handbook series linear algebra: Simultaneous iteration method for symmetric matrices. Numerische Mathematik, 16(3):205–223, 1970. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic).

Rutishauser:1970:SIM

[Rut70b] H. Rutishauser. Simultaneous iteration method for symmetric matrices. Numerische Mathematik, 16(3):205–223, December 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF02219773. Also in [WR71, pp. 284–302].

Rutishauser:1971:JMR

[Rut71a] H. Rutishauser. The Jacobi method for real symmetric matrices. In Wilkinson and Reinsch [WR71], pages 202–211. ISBN 0-387-05414-6, 3-540-05414-6, 3-662-38854-5, 3-662-39778-1. LCCN

QA251 .W67. URL http://link.springer.com/chapter/10.
1007/978-3-642-86940-2_12;   http://link.springer.com/
chapter/10.1007/978-3-662-39778-7_12.

Rutishauser:1971:SIM

[Rut71b] H. Rutishauser. Simultaneous iteration method for symmetric matrices. In Wilkinson and Reinsch [WR71], pages 284–302. ISBN 0-387-05414-6, 3-540-05414-6, 3-662-38854-5, 3-662-39778-1. LCCN

QA251 .W67. URL http://link.springer.com/chapter/10.
1007/978-3-642-86940-2_20;   http://link.springer.com/
chapter/10.1007/978-3-662-39778-7_20.

Rutishauser:1976:VNMa

[Rut76a] Heinz Rutishauser. Vorlesungen über numerische Mathematik. Band 1. Gleichungssysteme, Interpolation und Approximation. (German) [Lectures on numerical mathematics. Part 1. Equation systems, interpolation, and approximation], volume 50. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1976. ISBN 3-7643-0810-9. 164 pp. LCCN QA297 .R87. Herausgegeben von Martin Gutknecht unter Mitwirkung von Peter Henrici, Peter Läuchli und Hans-Rudolf Schwarz, Lehrbücher und Monographien aus dem Gebiete der exakten Wissenschaften. Mathematische Reihe.

Rutishauser:1976:VNMb

[Rut76b] Heinz Rutishauser. Vorlesungen über numerische Mathematik. Band 2. Differentialgleichungen und Eigenwertprobleme. (German) [Lectures on numerical mathematics. Volume 2 Differential equations and eigenvalue problems], volume 57. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1976. ISBN 3-7643-0850-8. 228 pp. Herausgegeben von Martin Gutknecht unter Mitwirkung von Peter Henrici, Peter Läuchli und Hans-Rudolf Schwarz, Lehrbücher und Monographien aus dem Gebiete der exakten Wissenschaften. Mathematische Reihe.

Rutishauser:1990:LNM

[Rut90]  Heinz Rutishauser.  Lectures on Numerical Mathematics.

Birkhäuser Boston Inc., Cambridge, MA, USA, 1990. ISBN 0-8176-3491-6, 1-4612-3468-9 (e-book), 1-4612-8035-4, 3-7643-3491-6. xv + 546 pp. LCCN QA297 .R8713 1990. URL http://www.

sam.math.ethz.ch/~mhg/pub/mhg-published/Rutishauser-LoNM.

pdf. Edited by Martin Gutknecht with the assistance of Peter Henrici, Peter Läuchli and Hans-Rudolf Schwarz, with a foreword by Gutknecht and a preface by Henrici, Läuchli and Schwarz, and translated from the German and with a preface by Walter Gautschi.

Rutishauser:2013:NAS

[Rut13] Hanna Rutishauser. Numerik, ALGOL und die Schweizer Hochalpen. Zur Arbeit an der Biografie von Heinz Rutishauser (1918–1970). (German) [Numerics, ALGOL and the Swiss High Alps. On the work on the biography of Heinz Rutishauser (1918–1970)]. Informatik Spektrum, 36(5):??, ???? 2013. CO-DEN INSKDW. ISSN 0170-6012 (print), 1432-122X (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00287-013-0730-z.

S:1978:BRM

[S.78] H. J. S. Book review: Martin Gutknecht, Peter Henrici, Peter Lauchli, and Hans-Rudolf Schwarz, Heinz Rutishauser: Vorlesungen Uber Numerische Mathematik Vols. 1 and 2. Mathematics of Computation, 32(142):651–652, April 1978. CODEN MCMPAF.

ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://

www.jstor.org/stable/pdf/2006176.pdf.

Saxer:1955:BGB

[SBRD^{+}55] W. Saxer, E. Baumann, E. Roth-Desmeules, K. Berger, H. Labhart, H. Rutishauser, and E. Völlm. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 6(3):252– 256, May 1955. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-

9039 (electronic). URL http://link.springer.com/article/
10.1007/BF01600645.

Schmid:1969:BRF

[Sch69] E. Schmid. Book review: F. L. Bauer and A. S. Householder and F. J. Olver and H. Rutishauser and K. Samelson and E. Stiefel, Handbook for Automatic computation, Vol. I. Die Grundlehren der mathematischen Wissenschaften in Einzeldarstellungen mit besonderer Berücksichtigung der Anwendungsgebiete, Band 135. Computing, 4(3):276, September 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmid:1970:BRF

[Sch70] E. Schmid. Book review: F. L. Bauer and A. S. Householder and F. W. J. Oliver and H. Rutishauser and K. Samelson and E. Stiefel (Hg.), Handbook for Automatic Computation, Vol. I, Part B (Die Grundlehren der mathematischen Wissenschaften in Einzeldarstellungen mit besonderer Berücksichtigung der Anwendungsgebiete: Band 137). Computing, 6(1–2):214, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schwarz:1981:EYC

[Sch81] H. R. Schwarz. The early years of computing in Switzerland. Annals of the History of Computing, 3(2):121–132, April/June 1981. CODEN AHCOE5. ISSN 0164-1239. URL http://dlib.

computer.org/an/books/an1981/pdf/a2121.pdf;   http://
www.computer.org/annals/an1981/a2121abs.htm.

Seewald:1982:QDA

[See82] Wolfgang Seewald. Quotienten-Differenzen-Algorithmus: Beweis der Regeln von Rutishauser. (German) [Quotient–difference algorithm: proof of Rutishauser’s rule]. Numerische Mathematik, 40 (1):93–98, November 1982. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic). URL http://link.springer.

com/article/10.1007/BF01459078.

Soriano:1988:RMM

[SF88] H. L. Soriano and F. Venâncio Filho. On Rutishauser’s modified method for computing the eigenvalues of real tridiagonal symmetric matrices. Computers and Structures, 29(1):151–154, 1988. CO-DEN CMSTCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Shield:1967:BGB

[SGR^{+}67] R. T. Shield, Th. Ginsburg, H. Rutishauser, K. Hutter, F. Weinberg, L. S. Dzung, W. Saxer, and P. Huguenin. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 18(2):298–300, March 1967. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http://link.springer.

com/article/10.1007/BF01596927.

Soriano:1982:RMM

[Sor82] Humberto Lima Soriano. Rutishauser’s modified method for computing the eigenvalues of symmetric matrices. Computer Methods in Applied Mechanics and Engineering, 35(3):255–269, 1982. CO-DEN CMMECC. ISSN 0045-7825 (print), 1879-2138 (electronic).

Stiefel:1961:RCI

[SR61] Edouard Stiefel and Heinz Rutishauser. Remarques concernant l’intégration numérique. (French) [Notes on numerical integration]. Comptes rendus de l’Académie des sciences, Paris, 252: 1899–1900, 1961. ISSN 0001-4036 (print), 2419-6304 (electronic).

Stutz:1953:BGB

[SRDH^{+}53] Th. Stutz, E. Roth-Desmeules, F. Held, F. Lüdi, and H. Rutishauser. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 4(2):163–166, March 1953. CODEN ZAM-PDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http:

//link.springer.com/article/10.1007/BF02067587.

Scherrer:1952:BGB

[SRP52] P. Scherrer, H. Rutischauser, and M. Plancherel. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 3(3):238–240, May 1952. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL https://link.springer.

com/article/10.1007/BF02008836.

Schwarz:1973:NAS

[SRS73] H. R. Schwarz, H. Rutishauser, and E. Stiefel. Numerical Analysis of Symmetric Matrices. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1973. ISBN 0-13-626556-1. xi + 276 pp. LCCN QA263 .S3313.

Sauer:1968:MHI

[SS68] R. Sauer and I. Szabó, editors. Mathematische Hilfsmittel des Ingenieurs. Teil III. (German) [Engineer’s mathematical tools. Part III], volume 141 of Die Grundlehren der mathematischen Wissenschaften. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. ISBN 3-642-95030-2, 3-642-95031-0. xix + 535 pp. LCCN QA76 .B3192. Herausgegeben von R. Sauer und I. Szabó. Unter Mitwirkung von H. Neuber, W. Nürnberg, K. Pöschl, E. Truckenbrodt und W. Zander.

Stiefel:1963:BGB

[SSG^{+}63] E. Stiefel, W. Saxer, W. Gerber, H. Rutishauser, P. Junod, U. Hochstrasser, W. Furrer, G. Epprecht, Ch. Clavuot, and W. Schumann. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 14(1):97–100, January 1963. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01601151.

Stetter:1969:BRH

[Ste69] H. J. Stetter. Book review: H. R. Schwarz, H. Rutishauser und E. Stiefel, Numerik symmetrischer Matrizen (Leitfaden der angewandten mathematik und Mechanik, Bd. 11). Computing, 4(4): 386, December 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stewart:1978:BRV

[Ste78] G. W. Stewart. Book review: Vorlesungen über numerische Mathematik, by Heinz Rutishauser. Bulletin of the American Mathematical Society, 84(4):660–664, July 1978. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Strang:1961:BRR

[Str61] W. Gilbert Strang. Book review: Refined iterative methods for computation of the solution and the eigenvalues of selfadjoint boundary value problems. By M. Engeli, Th. Ginsburg, H. Rutishauser, E. Stiefel. Birkhäuser Verlag, Basel, Stuttgart, 1959. 107 pp. $3.95. Quarterly of Applied Mathematics, 18(4): 374, 386, January 1961. CODEN QAMAAY. ISSN 0033-569X (print), 1552-4485 (electronic). URL https://www.jstor.org/

stable/43634808.

Symm:1981:EBC

[SW81] H. J. Symm and J. H. Wilkinson. Error bounds for computed invariant subspaces. Research Report 81-02, Seminar für Angewandte Mathematik, Eidgenössische Technische Hochschule, ETH-Zentrum CH-8092 Zurich, Switzerland, 1981. Lecture presented at a Symposium in honour of H. Rutishauser, October 15–17, 1980.

Todd:1952:BRP

[Tod52] John Todd. Book review: Programmgesteuerte digitale Rechengeräte (elektronische Rechenmaschinen), By H. Ruthishauser, A. Speiser, and E. Stiefel. (Mitteilungen aus dem Institut fur angewandte Mathematik, no. 2.) Basel, Birkhäuser, 1951. 102 pp. 8.50 Swiss fr. Bulletin of the American Mathematical Society, 58(2):278– 279, March 1952. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Thompson:2015:RAB

[TT15] Mary Clair Thompson and Tin-Yau Tam. Rutishauser’s LR algorithm and Bruhat iteration. Linear Multilinear Algebra, 63(10): 2061–2070, 2015. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

vandenHovedErtsenryck:2019:NIO

[van19] Gauthier Marie Claude Jean Thérèse Gérald van den Hove d’Ertsenryck. New Insights from Old Programs: The Structure of The First ALGOL 60 System. Ph.D. thesis, Centrum Wiskunde & Informatica (CWI), Informatics Institute, Faculty of Science, Universit of Amsterdam, Amsterdam, The Netherlands, Febru-

ary 2019. ???? + 302 pp. URL https://hdl.handle.net/
11245.1/8ecd6e9b-767d-44ab-84a7-8eed35822da5;   https:

//ir.cwi.nl/pub/28427/. Preface by Donald E. Knuth.

vonMuralt:1951:BGB

[vMNJ^{+}51] A. von Muralt, P. Niggli, M. Jeger, M. Strutt, Th. Erismann, H. Ziegler, W. Zünti, H. Rutishauser, A. P. Speiser, G. Trümpler, and W. Stutz. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 2(5):412–420, September 1951. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02579705.

Watkins:1990:RAS

[WE90] D. S. Watkins and L. Elsner. On Rutishauser’s approach to selfsimilar flows. SIAM Journal on Matrix Analysis and Applications, 11(2):301–311, April 1990. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Weiss:1988:BOP

[Wei88] Eric A. Weiss. Biographies: Oh, pioneers! Annals of the History of Computing, 10(4):348–361, October/December 1988. CODEN

AHCOE5. ISSN 0164-1239. URL http://dlib.computer.org/
an/books/an1988/pdf/a4348.pdf;   http://www.computer.
org/annals/an1988/a4348abs.htm.

Wilkinson:1971:LA

[WR71] James H. Wilkinson and Christian Reinsch, editors. Linear Algebra, volume II of Handbook for Automatic Computation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. ISBN 0-387-05414-6, 3-540-05414-6, 3-662-38854-5, 3-662-39778-1. viii + 439 pp. LCCN QA251 .W67. URL http://

www.springerlink.com/content/978-3-662-39778-7.

Weber:1963:BGB

[WRDF^{+}63] H. Weber, E. Roth-Desmeules, W. Furrer, R. Jost, H. R. Schwarz, D. Rivier, G. Epprecht, H. Rutishauser, P. Schmid, P. Läuchli, B. Federer, J. Ackeret, H. Schmid, W. Känzig, Hs. H. Günthard, and H. Ziegler. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 14(4):390–396, July 1963. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01603100.

Weber:1956:BGB

[WRDG^{+}56] H. Weber, E. Roth-Desmeules, Ed. Gerecke, E. Stiefel, H. Rutishauser, M. Jeger, W. Hoppe, H. Favre, J. Zeller, and A. A. Rusterholz. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 7(3):265–272, May 1956. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http://

link.springer.com/article/10.1007/BF02044475.

Ziegler:1962:BGB

[ZGR62] H. Ziegler, Ed. Gerecke, and H. Rutishauser. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 13 (2):199–200, March 1962. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http://link.springer.

com/article/10.1007/BF01594144.

Zielke:1981:VTR

[Zie81a] G. Zielke. Verallgemeinerungen einer Testmatrix von Rutishauser mit exakten Moore–Penrose-Inversen. (German) [Generalizations of a test matrix of Rutishauser with exact Moore–Penrose inverses]. Zeitschrift für Angewandte Mathematik und Mechanik, 61 (12):662–663, 1981. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Zielke:1981:GRT

[Zie81b] Gerhard Zielke. Generalizations of a Rutishauser test matrix with exact Moore–Penrose inverses. ACM SIGNUM Newsletter, 16(3): 7–8, September 1981. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Zuber:1966:BGB

[ZPS^{+}66] R. Zuber, P. Profos, Bruno Scarpellini, W. Saxer, P. J. Huber, A. Huber, H. R. Schwarz, G. Epprecht, A. A. Witt, Ch. Wehrli, H. Ziegler, H. Rutishauser, D. Rivier, D. B. McVean, F. Weinberg, and I. Zschokke-Gränacher. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 17(6): 794–800, November 1966. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http://link.springer.

com/article/10.1007/BF01595119.